### v1

# Insight-HXMT View of the BHC Swift J1727.8-1613 during its outburst in 2023

Abstract
The transient Galactic black hole candidate Swift J1727.8-1613 went through an outburst for the very first time that started in August 2023 and lasted for almost 6 months. We study the timing and spectral properties of this source using publicly available archival Insight-HXMT data for the first 10 observation IDs that last from MJD 60181 to 60198 with a total of 92 exposures for all three energy bands. We extracted the quasi-periodic oscillation properties by model fitting the power density spectrum and from those properties we designate that the QPOs are type-C in nature. We also conclude that the origin of the QPOs could be the shock instabilities, formed in the transonic accretion flow around black holes. The spectral analysis was performed using simultaneous data from the three on-board instruments LE, ME, and HE of HXMT in the broad energy band of $2-150$ keV. To achieve the best fit, we needed a combination of interstellar absorption, power-law, multi-color disk-blackbody continuum, gaussian emission/absorption, and power-law reflection by neutral material. From the spectral properties, we found that the source was in an intermediate state at the start of the analysis period and was making a transition toward the softer states. The shock (the boundary layer of the corona) moved inward in progressive days in accordance with the spectral nature. We found that the source is present in a high-inclination binary system. The hydrogen column density was found with an average value of $0.27_{-0.17}^{+0.08}\times 10^{22}$ cm-2.
X-rays: binaries – stars: black holes – stars: individual (Swift J1727.8-1613) – accretion, accretion disk – shock – radiation
journal: ApJ
1 Introduction
Black holes (BHs) are one of the end-products of massive stars. Stellar-mass black holes (SBHs) are one of the three types of presently detected BHs. They reside in binary systems with a companion star and most of their radiation comes out in the form of X-rays. Thus, they are called black hole X-ray binaries or BHXRBs. Depending on the mass of the companion star, they can be mainly of two types: low-mass X-ray binaries (LMXRBs) or high-mass X-ray binaries (HMXRBs). LMXRBs consist of type-A or later type of stars whereas the HMXRBs consist of giant O or B-type of stars (White et al. 1995; Tetarenko et al. 2016). Matter coming from the companion accumulates at a distance from the BH due to lack of viscosity, known as the pile-up radius (Chakrabarti et al. 2019; Chatterjee et al. 2022). Due to the accumulation of more matter, the temperature will rise. This gradual increment of temperature sets in instability at the pile-up, which therefore increases the viscosity. At some point in time when the accumulated matter will gain enough viscosity to push forward, accretion starts. On the other hand, irradiation of the accumulated matter at the outer boundary by the central object can also trigger mass accretion when the temperature of the accumulated matter crosses a critical limit (King & Ritter 1998; Mondal 2020). Thus, these systems make ideal laboratories for studying the physics of accretion. The BHXRBs mostly stay in a quiet dormant state when their flux stays at a very low detection level. Episodically, the radiation level increases by a very large factor and they become easily detectable. This is known as an outburst (Remillard & McClintock 2006). The X-ray luminosity increases by several orders of magnitude during an outburst when compared to a quiescence. Depending on the nature of the outburst, the sources are classified into two types: persistent and transient sources. Most of the BHXRBs are transient type of sources. They stay in the dormant state for most of their lifetimes. The luminosity becomes $\sim 10^{35}$ erg s-1 during episodic outbursting phase.
During the onset of an outburst, the flux in the light curve changes noticeably. According to nature, the outbursts are divided into two types (Debnath et al. 2010): fast rise slow decay (FRSD) and slow rise slow decay (SRSD). However, there are also some other classified outburst types, e.g., according to Zhang et al. (2019), outbursts can also be classified into glitch, reflare, multipeak, and mini-outburst. There are generally four designated spectral states during a complete BH outburst (Remillard & McClintock 2006). They are known as hard state (HS), hard intermediate state (HIMS), soft intermediate state (SIMS), and soft state (SS). Generally, when an outburst starts, it starts in the HS. Then it slowly moves toward the SS through the HIMS and SIMS (Belloni et al. 2005, 2011, and references within). After it reaches its SS the rising phase generally comes to an end and the outburst starts its decaying phase. In the decaying phase, it goes in the opposite direction and reaches the HS again. When a BH outburst goes through all these states, it is known as a type-I outburst. When the soft state is absent, it is known as a ‘failed’ or type-II outburst.
The radiation spectrum of a BH generally consists of two main components. One is the soft thermal multi-colour blackbody component, which is due to the blackbody radiation of the seed photons in the standard accretion disk (Shakura & Sunyaev 1973; Novikov & Thorne 1973). The other is the hard power-law component, which extends up to very high energies. The origin of this component is due to the inverse Comptonization of a fraction of seed photons from the disk, which are intercepted by a hot Compton cloud (Sunyaev & Titarchuk 1980, 1985; Haardt & Maraschi 1993; Zdziarski et al. 1993; Titarchuk 1994; Chakrabarti & Titarchuk 1995; Życki et al. 1999). When the outburst is in HS, there is very small contribution from the thermal component and mostly the non-thermal component dominates. In SS, the thermal component becomes dominant while in the intermediate states (HIMS and SIMS), the contribution from the two components stays comparable. other than these two components, other components could also be present in the radiation spectrum of a BH. In the case of a high soft state (HSS), there could be the presence of bulk motion Comptonization (BMC) which takes place due to the relativistic speed of matter when it reaches a very close distance to the BH (Blandford & Payne 1981a,b; Payne & Blandford 1981; Psaltis & Lamb 1997; Borozdin et al. 1999; Psaltis 2001;). It was been recently suggested by Chatterjee et al. (2023) that BMC was present in the radiation spectrum of the BHC MAXI J0637-430. The reprocessed emission plays an important role in the case of BHXRBs. Fraction of electromagnetic radiation can come back from the optically thin and geometrically thick Compton cloud to the optically thick and geometrically thin accretion disk due to the strong gravitational bending effect of the BH (Riaz et al. 2021). The illumination of the Comptonized photons from the Compton cloud to the accretion disk is known as returning or reprocessed radiation (George & Fabian 1991; Ross & Fabian 2005; Garcia & Kallman 2010). The reflection spectrum is a very important part of the whole radiation spectrum which helps determine many important properties of the source. The reflection radiation can produce some very specific features in the spectrum. One of which is the iron $K_{\alpha}$ emission line at $\sim 6.4$ keV. This happens when the returning radiation illuminates the disk and there is the fluorescence of ionized Fe. This line emission has a width of the order of $\sim 1~{}eV$ (Pozdniakov et al. 1979). However, for rotating black holes there are effects like Doppler shift and gravitational broadening which makes the line broadened and for rapidly rotating black holes this line can have two peaks due to red-shifting and blue-shifting of the component. This line emission is a very important aspect in estimating the spin of the BH. There could also be other features like Fe-K absorption edge and Compton hump. These features arise because of counterplay between two phenomena. Which feature will be prominent will be resulted in depending on which phenomenon is dominant. These are also energy-dependent. The interplay between photoelectric absorption and Compton scattering phenomena (Guilbert & Rees 1988; Lightman & White 1988; Fabian et al. 1989) will decide the spectral feature. At higher energies above $20$ keV, the scattering will be dominant resulting in a Compton hump, whereas, in low energies, there will be the dominance of photoelectric absorption and the spectrum drops away (Lightman & White 1988). As these features take place at a closer distance from the compact object (here a BH), we can notice broad and asymmetric reflection profiles in the spectrum due to distortion due to high gravitation. Proper modeling of the composite spectrum using different combinations of models can give good information about the physics that is going on there. In this regard, high signal-to-noise (s/n) data in the broadband range, e.g., Insight-HXMT (Zhang et al. 2020) can be very useful.
Apart from the spectral features and variabilities, there are temporal variabilities also. Time variability, which is proportional to the length scale of the photon-emitting region, is as important as X-ray spectroscopy to understand the physical properties of accretion flow dynamics in the region near to the BH. During an outburst, we notice that there are very small time-scale variabilities in the light curves, especially in the high-energy ones. If we make a Fourier transformation of the light curve, we notice that fast variability is present in the power density spectrum or PDS (van der Klis 1989). In the PDS, which is a variation of $P(\nu)$ with $\nu$, the variability is mainly aperiodic, which is a direct consequence that the accretion flow onto the Bh is turbulent itself. The aperiodic variability, which is the result of the combination of the variation of the source flux and the variation in Poisson statistics, contains both broadband noise and narrow features. The broadband noise is spread over a large frequency range, whereas the narrow feature is a power peak in narrow frequency ranges. The narrow peak is known as the quasi-periodic oscillation (QPO), which is one of the most important properties of a BH. There is a break frequency ($\nu_{b}$) that also characterizes the PDS (van der Klis 1994). The noise is modelled by the power-law function, whereas the QPOs are modelled by one or multiple Lorentzian profiles. In BHs, we see low-frequency QPOs (LFQPOs) mainly due to their geometrical origin. Depending on some properties like the frequency ($\nu$), $Q$-value ($=\nu/\delta\nu$, where $\delta\nu$ is the full width at half maximum or FWHM), RMS, etc., LFQPOs are classified into three different types: type A, B, and C (Casella et al. 2005). Although the clear origin of QPOs is still a matter of debate at present, several ideas were put forward as to why these features are being observed. Explanations like Lense-Thirring precession (Stella et al. 1999; Ingram et al. 2009), magneto-acoustic waves (Titarchuk et al. 1998), accretion-ejection instability (Tagger & Pellat 1999), shock oscillation model (Molteni et al. 1996; Chakrabarti et al. 2005, 2008, 2015) were put forward to explain the origin of LFQPOs in stellar-mass BHs. The shock oscillation model can give a simultaneous explanation of both the spectral and temporal properties, which makes it a very generally comprehensive model. High-resolution light curve data becomes very useful to study the temporal variability properties of BH. In this regard, Insight-HXMT is also very useful as well as studying the spectral properties.
Apart from these properties and features of a BH outburst, there are a few other estimations that give a direct correspondence between the spectral and temporal properties. A good correlation is observed between the two using just the light curve properties, e.g., the hardness ratio or HR, hardness intensity diagram or HID (Homan et al. 2001). Some model-fitted approaches are also important in this regard, e.g., accretion rate ratio or ARR, accretion rate ratio intensity diagram or ARRID (Mondal et al. 2014; Jana et al. 2016; Chatterjee et al. 2020). The RMS-intensity diagram or RID (Munoz-Darias et al. 2011), and hardness ration-intensity diagram or HRD (Belloni et al. 2005) are also very important estimates when it comes to correspondence between spectral and temporal properties. In different spectral states, all of these properties show different values, which are quite general for BH outbursts.
The BHC Swift J1727.8-1613 was discovered very recently on 24 August 2023 or MJD 60180 (Kennea & Swift Team 2023). The source was very bright as MAXI (Matsuoka et al. 2009) detected a flux of $7~{}Crab$ in the $2-20$ keV energy range. After that multi-wavelength follow-up observations were carried out by several ground-based and space-bourne telescope facilities (Baglio et al. 2023; Miller-Jones et al. 2023; Negoro et al. 2023b; O’Connor et al. 2023; Wang & Bellm 2023; Williams-Baldwin et al. 2023). The source is reported to have a distance of $2.7\pm 0.3kpc$ (MataSanchez et al. 2024) and a mass of $10\pm 2~{}M_{\odot}$ (Svoboda et al. 2024). The revelation of the presence of hydrogen and helium emission lines in the optical spectrum lead to its classification as an LMXRB (Castro-Tirado et al. 2023). The X-ray spectrum and other properties point to its nature as a BHXRB (Liu et al. 2023; Sunyaev et al. 2023). This was supported strongly after the detection of type C QPOs (Palmer & Parsotan 2023; Draghis et al. 2023; Bollemeijer et al. 2023; Mereminskiy et al. 2023) and flat radio spectrum (Miller-Jones et al. 2023; Bright et al. 2023). There is X-ray monitoring of the source with many satellites, e.g., NICER, NuSTAR, Insight-HXMT, IXPE, etc. Using IXPE data on September 7, 2023, Veledina et al. (2023) estimated its polarization with a polarization degree (PD) of $4.1\%\pm 0.2\%$ and PA of $32^{\circ}.2\pm 1^{\circ}.$3. There was a further polarization study of the source by Ingram et al. (2023). The presence of evolving LFQPOs at the high energy band of using AstroSat data was reported by Nandi et al. (2024). Using X-ray spectroscopy, there were reports about the spectral nature of the source. At the onset of the outburst, the spectral nature was consistent with a hard state with a photon index of power-law ($\Gamma$) $\sim 1.5-1.7$ (Liu et al. 2023). Then using the NICER observations, Bollemeijer et al. (2023) reported the softening of the spectrum with a substantial contribution of the soft blackbody component after 25 days of the start of the outburst. The presence of LFQPOs was reported by Draghis et al. (2023). Using NICER data on August 29 2023 (MJD 60185), Debnath et al. (2023) reported the presence of a soft time lag of $0.014\pm 0.001$ sec between energy bands of $3-10$ and $0.5-3$ keV at QPO frequency. With reflection spectroscopy, they found the inclination to be $\sim 87^{\circ}$.
Here is how this paper unfolds. In §1, we give an introduction. In §2, we describe the observation, data selection, reduction, and analysis. In §3, we portray the results. In §4, we make the discussion about the possible physical origin of our result. And finally, in §5, we make a summary of our results and draw some conclusions.
2 Observation, Data Selection, Reduction, and Analysis
This source is being extensively monitored by various X-ray satellites, e.g., NICER, NuSTAR, Insight-HXMT, IXPE, etc. For our analysis, we use China’s first dedicated X-ray satellite Insight-High X-ray Modulation Telescope (HXMT; Zhang et al. 2020). In the subsequent subsections, we discuss the data selection, reduction, and analysis, respectively.
2.1 Data Selection
This source was regularly monitored by the Insight-HXMT satellite. In the data archive, observations are available on-demand. At the time of the work, we found that there was a total of 34 observation IDs were available, although 3 of them were not downloadable. Out of those usable 31 observation IDs, we took the first 10 observation IDs. However, unlike other X-ray satellite data, each observation ID in HXMT has several different exposures. Also, some of the observation IDs continue for almost 2 to 3 days. We found that for these first 10 observation IDs, each of them has a minimum of 7 to a maximum of 18 exposures. We checked the analysis of all the exposures for the first 3 observation IDs and found that no significant change was noticed. Thus we did not fit every exposure for each observation ID. The details about the selected data are given in Table 1.
Table 1: List of Data used
Obs. Id.[1] Start Date[2] End Date[2] Start MJD[3] Stop MJD[3] Exp. (s)[4]
(1) (2) (3) (4) (5) (6)
P0614338001 2023-08-25 2023-08-27 60181.3410 60183.0597 148495
P0614338002 2023-08-27 2023-08-28 60183.0597 60184.2327 137343
P0614338003 2023-08-29 2023-08-31 60185.3075 60187.1589 159958
P0614338004 2023-08-31 2023-09-02 60187.1588 60189.0765 165684
P0614338005 2023-09-02 2023-09-04 60189.0765 60191.0601 171384
P0614338006 2023-09-04 2023-09-06 60191.0601 60193.4447 206034
P0614338008 2023-09-07 2023-09-08 60194.0350 60195.0926 91378
P0614338009 2023-09-08 2023-09-09 60195.0926 60196.0840 85659
P0614338010 2023-09-09 2023-09-10 60196.0840 60197.0754 85653
P0614338011 2023-09-10 2023-09-11 60197.0753 60198.0667 85650
[1] Column 1 lists each observation ID.
[2] In columns 2 and 3, we give the start and end date of each observation ID.
[3] In columns 4 and 5, we give the start and end MJD of each observation ID.
[4] Column 6 gives the exposure time of each observation ID.
2.2 Data Reduction
Insight-HXMT, China’s first astronomical X-ray satellite, was launched on 2017 June 15 (Zhang et al. 2014, 2020) to observe and study the properties of black holes, neutron stars (NSs), active galactic nuclei (AGNs), and other astronomical phenomena. The main scientific instrument is an array of 18 NaI/CSI phoswich scintillation detectors, each of which covers an effective area of 286 cm2. The satellite has three instruments or payloads: High Energy (HE; Liu et al. 2020), Medium Energy (ME; Chen et al. 2020), and Low Energy (LE; Cao et al. 2020). They have an effective area of 5100, 952, and 384 cm2 and cover an energy range of $20-250$, $5-30$, and $1-15$ keV, respectively.
After downloading on-demand level-1 data from the archive, we produced cleaned level-2 data for science analysis. This process of cleaning raw data was done in the following way. We first installed the HXMTDAS [^1] http://hxmt.org/index.php/usersp/dataan v2.05. Using this software, we run the hpipeline command using proper input and output directories. This pipeline performs several subsequent automatic commands for all three instruments. However, some suitable conditions are needed to be provided. Some conditions were set to get a good time interval, e.g., elevation angle $>10^{\circ}$; geomagnetic cutoff rigidity $>8~{}GeV$; pointing offset angle $<0.04^{\circ}$; $>600~{}s$ away from the South Atlantic Anomaly (SAA). All these commands altogether extract and clean the raw data and produce cleaned and analyzable science products. This is discussed in the HXMT manual [^2] http://hxmten.ihep.ac.cn/SoftDoc/501.jhtml in full details. The specific commands hespecgen, mespecgen, and lespecgen produce the spectra for HE. ME, and LE instruments. Whereas, the helcgen, melcgen, and lelcgen commands produce the light curve files for the three instruments. herspgen, merspgen, and lerspgen produce suitable response files. The background subtraction for both the timing and spectral data was performed using the commands hebkgmap, mebkgmap, and lebkgmap for instruments HE, ME, and LE, respectively. For $\chi^{2}$ fit-statistics in XSPEC, we group the spectrum using the grppha task of FTOOLS to a minimum of 30 counts per bin. We also set the time bin size to $0.01s$ to produce suitable light curves for generating PDS and looking for QPOs.
Using the cleaned spectra and light curves, we analyze the source properties, which are discussed in detail in the very next subsection.
2.3 Data Analysis
We perform timing and spectral analysis of the very first outburst during 2023 of the BHC Swift J1727.8-1613. We have downloaded the daily average light curve using the MAXI/GSC [^3] http://maxi.riken.jp/star_data/J1727-162/J1727-162.html public archive. Using light curves from all three modules of the HXMT, i.e., LE, ME, and HE, we have produced $0.01~{}s$ time-binned light curves first. Using those light curves, we have produced power density spectrum (PDS) by using the fast Fourier transformation (FFT) in the powspec task of the XRONOS package in the HEASoft software. Each observation’s data was subdivided into several intervals, each of which contains 8192 newbins. First, the PDS for each interval is generated and then they are averaged to make a resultant PDS. The PDS is normalized in a way that their integral will give the rms-squared fractional variability. White noise level is subtracted and a geometrical rebinning of -1.02 is used. For some data, we needed to use a geometrical rebinning of -1.05 to have a prominent QPO-like nature in the PDS. Using these steps, we searched for LFQPOs. We fit the QPO shape by the use of a Lorentzian model in powspec and extract QPO properties like frequency ($\nu_{qpo}$), full-width at half maximum or FWHM, normalization. For many observations, we have found the presence of single or multiple harmonics. We have also extracted their properties by fitting them with a Lorentzian model. We fitted the light curves of all the exposures of the listed 10 observations. We report them in the result section.
For spectral analysis, we have also used all three modules, i.e., LE, ME, and HE and fitted the broadband data in the $2-150$ keV energy band. Looking for the best fit, we have found that using the combination of disk blackbody, power-law, gaussian, and pexrav models, the data is best fitted. For some observations, we needed to add a gabs model to achieve the best fit. For interstellar absorption, we have used the tbabs model. Since we are simultaneously fitting all three modules, we have included a constant to normalize the three resultant fittings. The following two are our model combination: i) constant*tbabs*gabs*(diskbb + power-law + gaussian + pexrav) (for those for which gabs was required), ii) constant*tbabs*(diskbb + power-law + gaussian + pexrav) (for those for which gabs was not required. The spectral analysis was not done for all the exposures of the listed observations. Like the timing analysis, we did not spectrally analyze all the exposures, given in Table A. The spectrally analyzed exposures are marked with ‘*’ in Table A. The reason behind this is that, for timing analysis, we found variations of timing properties within a single day, whereas, for spectral analysis, the properties do not vary that much within such a short time. We report them in the result section.
We note that light curves for the LE instrument were not produced for some exposure IDs although we run the same pipeline for all three instruments. Thus we only represented those exposures MJDs in Table 3 for which all three light curves were available for analysis for uniformity.
3 Results
Using the timing and spectral analysis, we have studied the accretion flow properties of the very recent 2023 outburst of the BHC Swift J1727.8-1613. We studied the QPO properties of the source during the outburst as well as the spectral nature and radiation properties of the source using Insight-HXMT data. We broadly discuss the results in the two subsections below.
3.1 Temporal Properties
Here, we discuss the evolution of the flux and hardness ratio, and QPO properties during the first 10 observation IDs of the outburst.
Figure 1: Variation of the (a) MAXI/GSC $2-20$ keV count rate, (b) MAXI/GSC $2-6$ and $6-20$ keV count rates, and (c) hardness ratio with time. The HR is the ratio of the $6-20$ keV count rate to the $2-6$ keV count rate of the MAXI/GSC data.
3.1.1 Outburst Profile and Hardness Ratio
In Figure 1, we show the variation of the flux during the outburst. The outburst started roughly at around MJD 60180 (2023 August 24), when its flux came out of quiescence. The MAXI/GSC flux started to rise after this date, as we can notice from panel (a) of Figure 1. Within five days, the flux increased very rapidly and the total flux reached its peak on MJD 60185 (2023 August 29). Then the flux started to decrease very slowly. It decreased very slowly and gradually except for the period from MJD 60200 to 60225, where the flux showed a significant variation. In panel (b), we show the variations of the $2-6$ and $6-20$ keV energy bands count rates with time. We notice that, after MJD 60185, the hard flux (i.e., $6-20$ keV) decreased very slowly and gradually without showing any other variation, whereas, the soft $2-6$ keV flux showed a lot of variation just like the total $2-20$ keV flux. MJD 60185 is the peak for both the total $2-20$ keV and hard $6-20$ keV fluxes, whereas, the soft flux reached its peak much later at around MJD 60217. This suggests that the variation in the total count rate is mainly due to the variation of the soft component. After this period, the soft flux also decreased slowly to reach the end of the outburst and enter again into the quiescence phase.
In the panel (c), we show the variation of the hardness ratio, which is the ratio of the hard $6-20$ keV to the soft $2-6$ keV flux. At the start of the outburst, the HR was $\sim 0.5$. From this value, it first gradually decreased to MJD 60235 as the outburst progressed. After that, it became almost constant, although decreasing very slowly to $\sim 0.025$ at the end of our taken data of the outburst on MJD 60350. The hardness ratio is defined as $F_{H}/F_{S}$, where $F_{H}$ and $F_{S}$ are the hard and soft fluxes respectively. When this value is on the higher side, the spectral nature of the source is hard. This is the general case when an outburst is starting. When this value is low, there is dominance of soft flux over hard flux and the spectral nature would be soft. Without studying the detailed spectral analysis, HR provides a quick and rough idea about the spectral nature of an outburst. However, for this source, we notice that HR was a little higher at the onset of the outburst, where the hard flux was also comparable to the soft flux. However, after a few days, soft flux completely took over the hard flux, and the HR was also not high. Depending on the HR value, we propose that the source was already in the SIMS after the onset of the outburst. After MJD 60185, it made a transition into the SS. Since we are not analyzing the full duration of the outburst, it is not the scope of the paper to study the full spectral classification. We only spectrally define the source in the duration which falls into our analysis range.
3.1.2 Quasi Periodic Oscillations (QPOs)
As described in §2, we used the $0.01s$ time-binned light curves from all three bands (LE, ME, and HE) for our timing analysis. Using those light curves, we produced the power density spectrum (PDS) of those light curves to study QPOs. We find that QPO nature was present in all of the light curves in all three energy bands. In Figure 2, we show the variation of the QPO frequency ($\nu_{qpo}$) for the 10 different obs. IDs that we have used for our timing analysis. We first checked the first obs. ID P0614338001. It has several exposure IDs (given in Table 2). While we fit every exposure ID of this obs ID, we found that the QPO frequency was changing even within this small period. This is why we fitted all the light curve data of the exposures of the obs IDs that we have taken for analysis. We find that for every single obs ID, the $\nu_{qpo}$ showed significant variation. Figure 2(a), (b), (c) show the variations of the $\nu_{qpo}$ with time for 10 different obs IDs for LE, ME, and HE respectively.
Evolution of QPO frequency in different observation ID.s
Refer to captionRefer to caption
Figure 2: Evolution of QPO frequencies (Hz) with time. Here we have shown three different plots for (a) LE, (b) ME, and (c) HE instruments. Panels 1–10 in each of the three of these plots (a, b, c) represent the $\nu_{qpo}$ variation for the 10 observation IDs we used.
Figure 3: Evolution of QPO frequency with time during the whole period of analysis for (a) LE, (b) ME, and (c) HE.
In Figure 3, we show the evolution of the $\nu_{qpo}$ for the total duration of our analysis period. In (a), (b), and (c) panels the variations are for LE, ME, and HE respectively. We find that the QPO frequencies were $\sim 0.23\pm 0.008$, $0.24\pm 0.002$, and $0.24\pm 0.002$ Hz for LE, ME, and HE on the starting day of our analysis. After this day, the $\nu_{qpo}$ shows a rapid increase with time till $\sim$ MJD 60187 when $\nu_{qpo}$ was $\sim 1.2$Hz for all the three bands. After this date, it increased very slowly till MJD $\sim 60190$ when $\nu_{qpo}$ was $\sim 1.45$. Then it decreased for a very short period until it very rapidly increased to $1.5$Hz again in all three bands. After that, it showed varying nature till MJD $\sim 60197$ and then it finally increased again. This might be caused by the varying nature of the origin of these oscillations. We will discuss this in the next section.
Using the PDS fitting, we extracted some information (e.g., full-width at half maximum (FWHM), Normalization (LN)) of QPOs. We also extracted source and background count rates from the exposures. Using these estimates, we calculated the $Q$-value and RMS that represent the sharpness of the QPO and the fractional variability in the PDS, respectively. The values are listed in Table 3 in columns 5–7 ($Q$-value) and 8–10 (RMS) for LE, ME, and HE respectively.
Evolution of rms (%) with QPO frequency and time
Refer to captionRefer to caption
Figure 4: 3-D representation of the variation of $\nu_{qpo}$ and (%)RMS with time (MJD) for (a) LE, (b) ME, and (c) HE.
In Figure 4, we show a 3-D representation of the evolution of QPO RMS with corresponding frequency. In the LE band (Fig.4a), The RMS was higher at the start of the analysis period and it then gradually decreased. In the ME band (Fig. 4b), it increased from the onset of the analysis period for some time and then again decreased. HE band (Fig. 4c) shows a similar trend as of LE band. Since the QPO nature has already been reported in the recent paper by Yu et al. (2024), we would not put much insight into the classification of the QPO nature. However, by the values of both $Q$-factor and RMS, it can be said that the QPO is type-C in nature for all the exposure IDs that have been presented here.
3.2 Spectral Properties
Along with studying the timing properties, studying the spectral properties gives a clearer idea about the nature of the source. With the available Insight-HXMT data, we analyzed the source for a total of 32 exposure IDs. These exposure IDs are marked with a ‘*’ sign in the 1st column of Table 2. This source hasn’t been extensively analyzed spectrally. We perform an extensive study of this source’s spectral properties for the first few days after the onset of its outburst. We started our spectral analysis on MJD 60181.4. We have used LE+ME+HE in the $2-150$ keV energy band (LE in 2–10, ME in 10–35, and HE in 35–150 keV) for our spectral fitting for all these exposure IDs.
Model Fitted Insight-HXMT Spectra
Refer to captionRefer to caption
Figure 5: Model fitted spectrally analyzed unfolded spectra. (a) is or observation ID. P061433800501 (Exposure: P061433800501-20230902-01-01) for which the gabs model was necessary to achieve the best fit. (b) is for the observation ID. P061433800601 (exposure: P061433800601-20230904-01-01) for which the gabs model was not needed to achieve the best fit.
We first fitted the data with phenomenological diskbb and power-law models. We used the multiplicative tbabs model to account for the interstellar absorption. Our model read as constant*tbabs(diskbb + powerlaw) in XSPEC. The constant is taken to normalize between three different instruments in LE, ME, and HE. We found that the fit was not statistically acceptable. There was a clear sign of the Gaussian feature in the unfolded spectrum around $5-6$ keV. Also, there was a hump-like feature $\sim 20$ keV. Besides these features, the absorption feature was noticed below 2 keV. Thus, we refitted using the model combination as constant*tbabs*gabs(diskbb + power-law + ga + pexrav). Here, the gabs model takes care of the low energy absorption feature, while the ga and pexrav models take care of the contribution from the Gaussian and hump-like nature, as mentioned above. In Fig. 5a, we show the best-fitted unfolded spectrum using this model combination for the observation ID P061433800501 (Exposure ID: P061433800501-20230902-01-01), where $\chi^{2}/DOF\sim 1$. After some exposure, we found that the gabs model was not needed anymore to achieve the best fit. Then, the model combination in XSPEC reads as constant*tbabs(diskbb + power-law + ga + pexrav). In Figure 5b, we show an unfolded best-fitted spectrum using the later model combination. This is for the observation ID. P061433800601 (exposure ID: P061433800601-20230904-01-01), for which $\chi^{2}/DOF\sim 0.9$.
Model Fitted Insight-HXMT Spectra
Refer to captionRefer to caption
Figure 6: Variation of spectrally fitted properties (a) hydrogen column density ($n_{H}$), (b) inner-disk temperature ($T_{in}$), (c) diskbb normalization, (d) Photon Index ($\Gamma$), (e) reflection fraction (Rref), and (f) cutoff energy (Ecut) with time.
[…]
3.3 Correlation between spectral and timing properties
Photon Index ($\Gamma$) vs QPO frequency ($\nu_{qpo}$)
Refer to captionRefer to captionRefer to caption
Figure 7: Variation of photon index of power-law ($\Gamma$) with QPO frequency ($\nu_{qpo}$) for (a) LE, (b) ME, and (c) HE instruments. Here, $\Gamma$ is the spectral fitted index using combined LE+ME+HE data.
The QPO frequency could be produced from the oscillation of the Compton corona, which could also be the origin of the hard power-law photons that extend the spectrum to very high energies. They could be correlated to each other. It has been previously seen that the QPO frequency ($\nu_{qpo}$) and the photon index of power-law ($\Gamma$) show a positive correlation for some of the sources during their active outbursting phase (Vignarca et al. 2003; Shaposhnikov & Titarchuk 2009; Stiele et al. 2013). We also tried to find out if there are any correlations between these two properties. In Figure 7, we show the variations of the photon index with the QPO frequency for the three energy bands. Considering these points, we have found a positive correlation between $\Gamma$ and $\nu_{qpo}$. Using both the Pearson-Linear and Spearman-Rank correlation methods, we found correlation coefficients for the three energy bands as below.
Pearson-Linear Spearman-Rank
LE $0.65\pm 0.15$ $0.65\pm 0.19$
ME $0.65\pm 0.15$ $0.65\pm 0.19$
HE $0.64\pm 0.15$ $0.64\pm 0.19$
4 Discussions
The Galactic black hole Swift J1727.8-1613 was monitored daily by various X-ray satellites like NICER, NuSTAR, Insight-HXMT, etc. Multi-wavelength follow-up observations were also carried out for this source. We have taken HXMT data for both our spectral and timing analysis from MJD 60181.4 to 60198.0. The evolution of the light curve makes the outburst nature of the source a fast-rise slow decay type. After the onset of the outburst, it rose to its peak very fast within 5 days. Using the $0.01~{}s$ time-binned light curves from the three instruments of HXMT (LE, ME, and HE), we studied the source’s timing properties by analyzing the light curves. Using the spectra files from these three instruments we also studied the spectral properties of this source by analyzing combined LE+ME+HE spectra in the $2-150$ keV broad energy range.
4.1 Propagation of Shock
Figure 8: Variation of shock properties with time in accordance with $\nu_{qpo}$. In (a) the variation of $\nu_{qpo}$ with time and in (b) and (c) we show the variations of the shock location ($X_{s}$), and shock speed ($v$) with time.
Quasi-periodic oscillation is one of the most important and common phenomena for stellar mass black holes. For this newly discovered source, we analyzed light curves of a total of 342 light curves (114 for LE, ME, and HE each). However, in some of the exposures, light curves were not produced properly. There were a total of 92 exposures which commonly produced LE, ME, and HE light curves. We list those data in Table 3. Our analysis is final for 276 data. We found the presence of QPO in all of these light curves and they show rapid evolution in frequency in this short analysis period. The QPO frequency ($\nu_{qpo}$) showed an increase in frequency within even a single day. This is one of the advantages of using Insight-HXMT as it provides multiple exposures for one observation ID which helps to study the short-scale variations of such important physical properties of the source. The evolution of the QPO frequency has been discussed broadly in the result section. It showed an increasing trend from the start of our analysis. From the variation of QPO frequency, (%)RMS, and $Q$-factor, the nature of the QPO is designated to be type-C. Although this is a well-established topic in the literature, the origin of QPO is still a matter of debate. Many models try to explain the type-C QPO phenomenon, e.g., the Lense-Thirring effect (Stella & Vietri 1998), accretion-ejection instability (Tagger & Pellat 1999), oscillation of the disk (Titarchuk & Osherovich 2000), the misalignment between the spin and binary rotational axis (Ingram et al. 2009, and references therein).
We would like to focus on a physical scenario that the QPOs could have originated from the shock instabilities in advective flows around black holes. Black hole accretion is transonic in which multiple sonic points could exist (Chakrabarti & Titarchuk 1995). Depending on the flow properties, there is the possibility of the formation of standing shocks (Singh et al. 2022 and references therein). The matter from the companion need not necessarily be only Keplerian. There could be a supply of matter which has angular momentum distribution deviated from the Keplerian one. This is the sub-Keplerian component. Such a lower angular momentum component of infalling matter gets accreted in a free-fall timescale. Due to the tug of war between the pressure of gravitational force and centrifugal force, this matter could virtually stop from falling onto the BH at some distance from the BH and undergo shock transition, creating a post-shock region. This approach could very well describe the spectral and temporal properties of BH, as it has been shown previously in both theoretical (Chakrabarti & Titarchuk 1995) and observational (Debnath et al. 2014; Mondal et al. 2014, Chatterjee et al. 2020, 2021, 2023) studies. The change in size of such post-shock region can describe the spectro-temporal properties during an outburst. According to Molteni et al. (1996), the oscillation in the shock region could have originated due to the presence of cooling. This was later shown in simulation works by Garain et al. (2014). The oscillation of the shock region could have originated due to the resonance condition between the cooling timescale due to the Comptonization process and the infall timescale in the coronal region (see Chakrabarti et al. 2015). Additionally, it has also been observed by the authors that once the QPO sets in, it gets locked for a week or more depending on the above condition. From the expressions of infall and cooling timescales, we get that the QPO frequency ($\nu_{qpo}$) could be given as (Molteni et al. 1996; Chakrabarti & Manickam 2000; Chakrabarti et al. 2004),
$\nu_{qpo}=\frac{c^{3}}{2GM_{BH}}\frac{1}{RX_{s}(X_{s}-1)^{1/2}}Hz,$ (1)
where $c$ and $G$ are the speed of light and gravitational constant respectively. $M_{BH}$ represents the mass of the black hole, $X_{s}$ is the distance of the shock, or called shock location. $X_{s}$ is in the units of Schwarzschild radius ($r_{s}=2GM_{BH}/c^{2}$). $R$ is the ratio of matter densities in post-shock to pre-shock regions ($\rho_{+}/\rho_{-}$) and is known as the compression ratio. From our timing analysis, we already extracted the information about QPO frequency ($\nu_{qpo}$). Using the above relation, we estimated the shock location during the outburst. According to Chakrabarti et al. (2005), $X_{s}$ could be located anywhere from 10 to 1000 $r_{s}$. The shock stays very far away $\sim 1000~{}r_{s}$ when the spectral nature of the ongoing outburst is hard and becomes smaller gradually in the process of cooling. Using the relation (1), we found that the shock was far at the start of the analysis period (Fig. 8b). It then rapidly decreased as the QPO frequency increased rapidly during the initial few days. Close to MJD 60185, when the $\nu_{qpo}$ stopped its rapid increment, and the shock location started to decrease very slowly. The values of the shock location along with QPO frequency are given in Table 3. As equation (1) suggests, as the shock was bigger at the start of the outburst, the QPO frequency was small. As it decreased rapidly due to cooling, the $\nu_{qpo}$ increased rapidly. Then when the cooling was less efficient, the size of the shock changed very slowly, although decreasing. It made the $\nu_{qpo}$ increase with time but not so rapidly as before.
If the time of the first exposure ID was taken to be 0, and the shock location as $X_{s0}$ and after time $t$, if the shock is at $X_{s}$, then the velocity of the movement of the shock would be given by the equation (Chakrabarti et al. 2005)
[…]
Using this relation, we estimated the velocity of shock. This is given in Fig. 8c. We find that the shock velocity has varied within this period. In the beginning, it attained higher values when the $X_{s}$ was also larger, however decreasing rapidly. It slowed down as the shock moved inwards, along with some fluctuations. This could be due to the limitation of assuming constant velocity of the movement of the shock.
4.2 Evolution of the physical configuration and spectral nature
From the spectral properties, we have seen that the $T_{in}$ was first low and it gradually increased. The exact opposite trend was observed for the Normalization of the diskbb model. The norm of diskbb model varies as
[…]
where, $R_{in}$, $D_{10}$ are the inner-disk radius of the accretion disk and the distance of the source in $10$ kpc unit. $i$ is the inclination angle of the accretion disk to the observer. Considering, $D_{10}$ and $\theta$ do not change during an outburst, the relation could be shortened like $R_{in}\sim N_{diskbb}^{1/2}$. Thus as the $N_{diskbb}$ decreased, it means the inner edge of the disk moved closer to the black hole. This is also supported by the fact that $T_{in}$ increased as the disk came closer.
This also supports the properties we showed from the timing analysis. The shock forms at a distance ($X_{s}$) when the sub-Keplerian matter is halted due to the counterbalance between the pressures of gravitational force and the centrifugal force. This forms the hot Compton cloud, as it becomes hot and puffed up in time. The hot Compton cloud will intercept some of those disk photons and there will be scattering in the process of inverse-Comptonization that will increase the energy of those intercepted photons and will emit them at higher energies. In this way, the Compton cloud will be cooled due to energy loss. Due to this cooling, the Compton cloud will gradually shrink in size. As it continues shrinking in size, the disk will continue to move inwards. As we have seen from the timing properties, the shock location was larger at the start of the outburst. The inner-disk radius was also higher as $R_{in}\sim N_{diskbb}^{1/2}$ and $N_{diskbb}$ was higher. Then cooling starts as the disk starts to approach inwards. In this way, the Compton cloud was losing energy by up-scattering and as the cooling continued, the $X_{s}$ and $R_{in}$ were both moving closer to the black hole.
Looking at the spectral parameters, and also the location of the shock, we designate that the source was already in an intermediate state at the start of our analysis period. As the outburst progressed, it was transitioning towards softer states.
From the values of the parameter $\cos(i)$ of the pexrav model, we find that the inclination of the disk was $>60^{\circ}$, which makes this source a high inclination binary system.
The hydrogen column density ($N_{H}$) showed a variation, which is common for Galactic black hole binaries unless the system is heavily obscured, which has recently been found for the newly discovered BH Swift J151857.0-572147 (Mondal et al. 2024). The average $N_{H}$ has an average value of $0.27_{-0.17}^{+0.08}\times 10^{22}$ cm-2.
5 Summary and Conclusions
We have studied the timing and spectral properties of the BHC Swift J1727.8-1613 during its recent outburst in 2023. We show the evolution of the light curve along with its hardness ratio for the entire duration of the outburst using archival MAXI/GSC data. Studying the entire evolution of the outburst was not the scope of this paper. Using the Insight-HXMT data, we chose the first ten observation IDs for our analysis. We use all the exposures from those observation IDs and selective exposures for timing and spectral analysis, respectively. Using $0.01~{}s$ time-binned light curves from all three instruments of HXMT, i.e., LE, ME, and HE, we studied the QPO properties by producing a power density spectrum. The QPO properties were extracted with the use of Lorentzian model. For spectral analysis, we use LE + ME + HE spectrum files in the broad $2-150$ keV energy band. We found that the models i) constant*tbabs*gabs*(diskbb + power-law + gaussian + pexrav) (for those for which gabs was required), ii) constant*tbabs*(diskbb + power-law + gaussian + pexrav) fits the spectra best. From our analysis, we conclude that:
(i) Quasi-periodic oscillations were present during the entire period of our analysis, which showed evolution in QPO frequency.
(ii) The $Q$-factor and $RMS$ help designate the nature of the QPOs as type-C.
(iii) QPOs originated from the shock instabilities that was formed by matter having an angular momentum distribution deviated from the Keplerian distribution. As the shock moved inwards, the QPO frequency was observed to increase.
(iv) With the shock moving inwards, the spectral nature of the source was transitioning towards the softer states, after it started in the intermediate state at the onset of the analysis period.
(v) The spectral and temporal properties correlate to each other which makes our claim stronger about the origin of QPOs.
(vi) This source is in a high inclination binary system.
(vii) The hydrogen column density varied in the range of $0.10-0.35\times 10^{22}$ cm-2, which is quite common for interstellar absorption of Galactic black holes.
6 Data Availability
This work has made use of public data from several satellite/instrument archives and has made use of software from the HEASARC, which is developed and monitored by the Astrophysics Science Division at NASA/GSFC and the High Energy Astrophysics Division of the Smithsonian Astrophysical Observatory. This work made use of the data from the Insight-HXMT mission, a project funded by the China National Space Administration (CNSA) and the Chinese Academy of Sciences (CAS).
7 Acknowledgements
KC acknowledges support from the SWIFAR postdoctoral fund of Yunnan University. SM acknowledges the Ramanujan Fellowship (# RJF/2020/000113) by SERB-DST, Govt. of India for this research. CBS is supported by the National Natural Science Foundation of China under grant no. 12073021.

### v2

@@
 We extracted the quasi-periodic oscillation properties by model fitting the power density spectrum and from those properties we designate that the QPOs are type-C in nature.
-We also conclude that the origin of the QPOs could be the shock instabilities, formed in the transonic accretion flow around black holes.
-The spectral analysis was performed using simultaneous data from the three on-board instruments LE, ME, and HE of HXMT in the broad energy band of $2-150$ keV.
-To achieve the best fit, we needed a combination of interstellar absorption, power-law, multi-color disk-blackbody continuum, gaussian emission/absorption, and power-law reflection by neutral material.
+We also conclude that the origin of the QPOs could be the shock instabilities in the transonic advective accretion flows around black holes.
+The spectral analysis was performed using simultaneous data from the three on-board instruments LE, ME, and HE of Insight-HXMT in the broad energy band of $2-150$ keV.
+To achieve the best fit, spectral fitting required a combination of models e.g.
+interstellar absorption, power-law, multi-color disk-blackbody continuum, gaussian emission/absorption, and reflection by neutral material.
 From the spectral properties, we found that the source was in an intermediate state at the start of the analysis period and was making a transition toward the softer states.
-The shock (the boundary layer of the corona) moved inward in progressive days in accordance with the spectral nature.
-We found that the source is present in a high-inclination binary system.
-The hydrogen column density was found with an average value of $0.27_{-0.17}^{+0.08}\times 10^{22}$ cm-2.
+The inner edge of the accretion disk moved inward in progressive days following the spectral nature.
+We found that the source has a high inclination.
+The average hydrogen column density estimated from the model fitting is $0.27_{-0.17}^{+0.08}\times 10^{22}$ cm-2.
 X-rays: binaries – stars: black holes – stars: individual (Swift J1727.8-1613) – accretion, accretion disk – shock – radiation
-journal: ApJ
+journal: AAS
 1 Introduction
-Black holes (BHs) are one of the end-products of massive stars.
-Stellar-mass black holes (SBHs) are one of the three types of presently detected BHs.
-They reside in binary systems with a companion star and most of their radiation comes out in the form of X-rays.
-Thus, they are called black hole X-ray binaries or BHXRBs.
+Stellar-mass black holes (SBHs) are one of the end products of the death of massive stars.
+They reside in binary systems where a companion star supplies matter and during accretion they emit radiation.
+A significant amount of their radiation comes out in the form of X-rays, therefore, they are called black hole X-ray binaries (BHXRBs).
 Depending on the mass of the companion star, they can be mainly of two types: low-mass X-ray binaries (LMXRBs) or high-mass X-ray binaries (HMXRBs).
-LMXRBs consist of type-A or later type of stars whereas the HMXRBs consist of giant O or B-type of stars (White et al.
+The LMXRBs consist of type-A or later type of stars, whereas the HMXRBs consist of giant O or B-type of stars (White et al.
 1995; Tetarenko et al.
 2016).
-Matter coming from the companion accumulates at a distance from the BH due to lack of viscosity, known as the pile-up radius (Chakrabarti et al.
+During the transfer of mass from the companion, matter accumulates at a distance from the BH due to insufficient viscosity to drive the matter inward, which is called the pile-up radius (Chakrabarti et al.
 2019; Chatterjee et al.
@@
 Due to the accumulation of more matter, the temperature will rise.
-This gradual increment of temperature sets in instability at the pile-up, which therefore increases the viscosity.
-At some point in time when the accumulated matter will gain enough viscosity to push forward, accretion starts.
+The gradual increase in temperature develops instability at the pile-up radius, which further increases the viscosity.
+At some point in time when the accumulated matter gains enough viscosity to push forward, accretion starts.
 On the other hand, irradiation of the accumulated matter at the outer boundary by the central object can also trigger mass accretion when the temperature of the accumulated matter crosses a critical limit (King & Ritter 1998; Mondal 2020).
 Thus, these systems make ideal laboratories for studying the physics of accretion.
-The BHXRBs mostly stay in a quiet dormant state when their flux stays at a very low detection level.
-Episodically, the radiation level increases by a very large factor and they become easily detectable.
-This is known as an outburst (Remillard & McClintock 2006).
+The BHXRBs mostly stay in a quiet dormant state when their flux remains at a very low detection level.
+Episodically, the radiation level increases by a very large factor and they become easily detectable, which is known as an outburst (Remillard & McClintock 2006).
 The X-ray luminosity increases by several orders of magnitude during an outburst when compared to a quiescence.
@@
 During the onset of an outburst, the flux in the light curve changes noticeably.
-According to nature, the outbursts are divided into two types (Debnath et al.
+Based on its nature, the outbursts are divided into two types (Debnath et al.
 2010): fast rise slow decay (FRSD) and slow rise slow decay (SRSD).
@@
 There are generally four designated spectral states during a complete BH outburst (Remillard & McClintock 2006).
-They are known as hard state (HS), hard intermediate state (HIMS), soft intermediate state (SIMS), and soft state (SS).
+They are known as the hard state (HS), hard intermediate state (HIMS), soft intermediate state (SIMS), and soft state (SS).
 Generally, when an outburst starts, it starts in the HS.
@@
 2005, 2011, and references within).
-After it reaches its SS the rising phase generally comes to an end and the outburst starts its decaying phase.
-In the decaying phase, it goes in the opposite direction and reaches the HS again.
+After it reaches its SS, the rising phase generally comes to an end and the outburst starts its decaying phase.
+In the decaying phase, it goes in the opposite direction and reaches the declining HS.
 When a BH outburst goes through all these states, it is known as a type-I outburst.
 When the soft state is absent, it is known as a ‘failed’ or type-II outburst.
-The radiation spectrum of a BH generally consists of two main components.
+The radiation spectrum of a BHXRB generally consists of two main components.
 One is the soft thermal multi-colour blackbody component, which is due to the blackbody radiation of the seed photons in the standard accretion disk (Shakura & Sunyaev 1973; Novikov & Thorne 1973).
-The other is the hard power-law component, which extends up to very high energies.
-The origin of this component is due to the inverse Comptonization of a fraction of seed photons from the disk, which are intercepted by a hot Compton cloud (Sunyaev & Titarchuk 1980, 1985; Haardt & Maraschi 1993; Zdziarski et al.
+The other is the hard power-law component, which can extend up to very high energies.
+The origin of this component is due to the inverse Comptonization of a fraction of seed photons from the disk, which is intercepted by a hot Compton cloud (Sunyaev & Titarchuk 1980, 1985; Haardt & Maraschi 1993; Zdziarski et al.
 1993; Titarchuk 1994; Chakrabarti & Titarchuk 1995; Życki et al.
 1999).
-When the outburst is in HS, there is very small contribution from the thermal component and mostly the non-thermal component dominates.
-In SS, the thermal component becomes dominant while in the intermediate states (HIMS and SIMS), the contribution from the two components stays comparable.
-other than these two components, other components could also be present in the radiation spectrum of a BH.
-In the case of a high soft state (HSS), there could be the presence of bulk motion Comptonization (BMC) which takes place due to the relativistic speed of matter when it reaches a very close distance to the BH (Blandford & Payne 1981a,b; Payne & Blandford 1981; Psaltis & Lamb 1997; Borozdin et al.
-1999; Psaltis 2001;).
-It was been recently suggested by Chatterjee et al.
-(2023) that BMC was present in the radiation spectrum of the BHC MAXI J0637-430.
-The reprocessed emission plays an important role in the case of BHXRBs.
-Fraction of electromagnetic radiation can come back from the optically thin and geometrically thick Compton cloud to the optically thick and geometrically thin accretion disk due to the strong gravitational bending effect of the BH (Riaz et al.
-2021).
-The illumination of the Comptonized photons from the Compton cloud to the accretion disk is known as returning or reprocessed radiation (George & Fabian 1991; Ross & Fabian 2005; Garcia & Kallman 2010).
-The reflection spectrum is a very important part of the whole radiation spectrum which helps determine many important properties of the source.
-The reflection radiation can produce some very specific features in the spectrum.
-One of which is the iron $K_{\alpha}$ emission line at $\sim 6.4$ keV.
-This happens when the returning radiation illuminates the disk and there is the fluorescence of ionized Fe.
-This line emission has a width of the order of $\sim 1~{}eV$ (Pozdniakov et al.
-1979).
-However, for rotating black holes there are effects like Doppler shift and gravitational broadening which makes the line broadened and for rapidly rotating black holes this line can have two peaks due to red-shifting and blue-shifting of the component.
-This line emission is a very important aspect in estimating the spin of the BH.
-There could also be other features like Fe-K absorption edge and Compton hump.
-These features arise because of counterplay between two phenomena.
-Which feature will be prominent will be resulted in depending on which phenomenon is dominant.
-These are also energy-dependent.
-The interplay between photoelectric absorption and Compton scattering phenomena (Guilbert & Rees 1988; Lightman & White 1988; Fabian et al.
-1989) will decide the spectral feature.
-At higher energies above $20$ keV, the scattering will be dominant resulting in a Compton hump, whereas, in low energies, there will be the dominance of photoelectric absorption and the spectrum drops away (Lightman & White 1988).
-As these features take place at a closer distance from the compact object (here a BH), we can notice broad and asymmetric reflection profiles in the spectrum due to distortion due to high gravitation.
-Proper modeling of the composite spectrum using different combinations of models can give good information about the physics that is going on there.
-In this regard, high signal-to-noise (s/n) data in the broadband range, e.g., Insight-HXMT (Zhang et al.
+When the outburst is in HS, there is a very small contribution from the thermal blackbody component and mostly the (non)thermal powerlaw component dominates.
+In SS, the thermal blackbody component becomes dominant while in the intermediate states (HIMS and SIMS), the contribution from the two components stays comparable.
+Other than these two components, in the case of a high soft state (HSS), there could be the presence of bulk motion Comptonization (BMC) which takes place due to the relativistic speed of matter when it reaches a closer to the BH (Blandford & Payne 1981a,b; Payne & Blandford 1981; Chakrabarti & Titarchuk 1995; Psaltis & Lamb 1997; Borozdin et al.
+1999; Psaltis 2001).
+It was recently observed by Chatterjee et al.
+(2023) that BMC was present in the radiation spectrum of the black hole candidate (BHC) MAXI J0637-430.
+In addition, reprocessing of the Comptonized photons from the Compton cloud by the accretion disk is known as reprocessed radiation (George & Fabian 1991; Ross & Fabian 2005; Garcia & Kallman 2010), which generates reflection spectra.
+As a result, the Fe $K\alpha$ emission line at $\sim 6.4$ keV and a reflection hump above 20 keV may originate.
+Therefore, the presence of complex features in the spectrum in BHXRBs makes them important to study to understand their origin.
+This can be achieved by fitting broadband data using multicomponent models.
+In this regard, high-resolution broadband data of Insight-HXMT (Zhang et al.
 2020) can be very useful.
-Apart from the spectral features and variabilities, there are temporal variabilities also.
-Time variability, which is proportional to the length scale of the photon-emitting region, is as important as X-ray spectroscopy to understand the physical properties of accretion flow dynamics in the region near to the BH.
-During an outburst, we notice that there are very small time-scale variabilities in the light curves, especially in the high-energy ones.
-If we make a Fourier transformation of the light curve, we notice that fast variability is present in the power density spectrum or PDS (van der Klis 1989).
-In the PDS, which is a variation of $P(\nu)$ with $\nu$, the variability is mainly aperiodic, which is a direct consequence that the accretion flow onto the Bh is turbulent itself.
-The aperiodic variability, which is the result of the combination of the variation of the source flux and the variation in Poisson statistics, contains both broadband noise and narrow features.
-The broadband noise is spread over a large frequency range, whereas the narrow feature is a power peak in narrow frequency ranges.
-The narrow peak is known as the quasi-periodic oscillation (QPO), which is one of the most important properties of a BH.
-There is a break frequency ($\nu_{b}$) that also characterizes the PDS (van der Klis 1994).
-The noise is modelled by the power-law function, whereas the QPOs are modelled by one or multiple Lorentzian profiles.
-In BHs, we see low-frequency QPOs (LFQPOs) mainly due to their geometrical origin.
-Depending on some properties like the frequency ($\nu$), $Q$-value ($=\nu/\delta\nu$, where $\delta\nu$ is the full width at half maximum or FWHM), RMS, etc., LFQPOs are classified into three different types: type A, B, and C (Casella et al.
+Besides rich spectral features and variabilities, temporal properties are equally important to understanding the dynamics of the accreting gas around the BHs.
+During an outburst, it has been observed that the light curves show very small timescale variabilities, especially in the high-energy bands.
+The Fourier transformation of the light curve imprints such variabilities as both broadband noise and narrow features in the power density spectrum or PDS (van der Klis 1989).
+The broadband noise is spread over a large frequency range modeled by power-law function.
+The narrow feature is a power peak in narrow frequency ranges, known as the quasi-periodic oscillation (QPO), modeled by one or multiple Lorentzian profiles.
+In BHXRBs, the low-frequency QPOs (LFQPOs) are often observed mainly due to their geometrical origin.
+Depending on their properties like the frequency ($\nu$), $Q$-value ($=\nu/\delta\nu$, where $\delta\nu$ is the full width at half maximum or FWHM), (%)RMS, etc., LFQPOs are classified into three different types: type A, B, and C (Casella et al.
 2005).
-Although the clear origin of QPOs is still a matter of debate at present, several ideas were put forward as to why these features are being observed.
-Explanations like Lense-Thirring precession (Stella et al.
+Although their origin is still a topic of debate, several ideas were put forward e.g., Lense-Thirring precession (Stella et al.
 1999; Ingram et al.
@@
 1996; Chakrabarti et al.
-2005, 2008, 2015) were put forward to explain the origin of LFQPOs in stellar-mass BHs.
-The shock oscillation model can give a simultaneous explanation of both the spectral and temporal properties, which makes it a very generally comprehensive model.
-High-resolution light curve data becomes very useful to study the temporal variability properties of BH.
-In this regard, Insight-HXMT is also very useful as well as studying the spectral properties.
-Apart from these properties and features of a BH outburst, there are a few other estimations that give a direct correspondence between the spectral and temporal properties.
-A good correlation is observed between the two using just the light curve properties, e.g., the hardness ratio or HR, hardness intensity diagram or HID (Homan et al.
+2005, 2008, 2015).
+The shock oscillation model can explain both the spectral and temporal properties simultaneously, which makes it a more comprehensive model.
+As both the properties are coming from the same system, and changes in spectral states can also dictate the types of QPOs, it is believed that both are interlinked.
+A good correlation is observed between the two using just the light curve properties, e.g., the hardness ratio or HR, hardness intensity diagram, or HID (Homan et al.
 2001).
-Some model-fitted approaches are also important in this regard, e.g., accretion rate ratio or ARR, accretion rate ratio intensity diagram or ARRID (Mondal et al.
+Some model-fitted approaches also showed that the changes in both properties are due to the interplay between different mass accretion rates or variation in viscosity, can be understood through accretion rate ratio intensity diagram or ARRID (Mondal et al.
 2014; Jana et al.
@@
 The RMS-intensity diagram or RID (Munoz-Darias et al.
-2011), and hardness ration-intensity diagram or HRD (Belloni et al.
-2005) are also very important estimates when it comes to correspondence between spectral and temporal properties.
-In different spectral states, all of these properties show different values, which are quite general for BH outbursts.
+2011), and hardness ratio-intensity diagram or HRD (Belloni et al.
+2005) can also make a link between spectral and temporal properties from pure observational ground.
 The BHC Swift J1727.8-1613 was discovered very recently on 24 August 2023 or MJD 60180 (Kennea & Swift Team 2023).
-The source was very bright as MAXI (Matsuoka et al.
-2009) detected a flux of $7~{}Crab$ in the $2-20$ keV energy range.
-After that multi-wavelength follow-up observations were carried out by several ground-based and space-bourne telescope facilities (Baglio et al.
+MAXI (Matsuoka et al.
+2009) detected the source to be very bright with a flux of $7$ Crab in the $2-20$ keV energy range.
+Since then multi-wavelength follow-up observations were carried out by several ground-based and space-bourne telescope facilities (Baglio et al.
 2023; Miller-Jones et al.
@@
 2023).
-The source is reported to have a distance of $2.7\pm 0.3kpc$ (MataSanchez et al.
+The source is reported to be located at a distance of $2.7\pm 0.3$ kpc (MataSanchez et al.
 2024) and a mass of $10\pm 2~{}M_{\odot}$ (Svoboda et al.
 2024).
-The revelation of the presence of hydrogen and helium emission lines in the optical spectrum lead to its classification as an LMXRB (Castro-Tirado et al.
+The revelation of the presence of hydrogen and helium emission lines in the optical spectrum led to its classification as an LMXRB (Castro-Tirado et al.
 2023).
@@
 2023; Mereminskiy et al.
-2023) and flat radio spectrum (Miller-Jones et al.
+2023) and a flat radio spectrum (Miller-Jones et al.
 2023; Bright et al.
 2023).
-There is X-ray monitoring of the source with many satellites, e.g., NICER, NuSTAR, Insight-HXMT, IXPE, etc.
+There is X-ray monitoring of the source by NICER, NuSTAR, Insight-HXMT, IXPE, etc.
 Using IXPE data on September 7, 2023, Veledina et al.
-(2023) estimated its polarization with a polarization degree (PD) of $4.1\%\pm 0.2\%$ and PA of $32^{\circ}.2\pm 1^{\circ}.$3.
+(2023) estimated its polarization with a polarization degree (PD) of $4.1\%\pm 0.2\%$ and polarization angle (PA) of $32^{\circ}.2\pm 1^{\circ}.$3.
 There was a further polarization study of the source by Ingram et al.
 (2023).
-The presence of evolving LFQPOs at the high energy band of using AstroSat data was reported by Nandi et al.
+At the starting of the outburst, the spectral nature was consistent with a hard state with a power-law photon index ($\Gamma$) $\sim 1.5-1.7$ (Liu et al.
+2023).
+Later, using the NICER observations, Bollemeijer et al.
+(2023) reported the softening of the spectrum with a substantial contribution of the soft blackbody component after 25 days from the starting of the outburst.
+The presence of LFQPOs was reported by Draghis et al.
+(2023) and the evolution of LFQPOs at the high energy band using AstroSat data was reported by Nandi et al.
 (2024).
-Using X-ray spectroscopy, there were reports about the spectral nature of the source.
-At the onset of the outburst, the spectral nature was consistent with a hard state with a photon index of power-law ($\Gamma$) $\sim 1.5-1.7$ (Liu et al.
-2023).
-Then using the NICER observations, Bollemeijer et al.
-(2023) reported the softening of the spectrum with a substantial contribution of the soft blackbody component after 25 days of the start of the outburst.
-The presence of LFQPOs was reported by Draghis et al.
-(2023).
-Using NICER data on August 29 2023 (MJD 60185), Debnath et al.
-(2023) reported the presence of a soft time lag of $0.014\pm 0.001$ sec between energy bands of $3-10$ and $0.5-3$ keV at QPO frequency.
-With reflection spectroscopy, they found the inclination to be $\sim 87^{\circ}$.
-Here is how this paper unfolds.
-In §1, we give an introduction.
-In §2, we describe the observation, data selection, reduction, and analysis.
-In §3, we portray the results.
-In §4, we make the discussion about the possible physical origin of our result.
-And finally, in §5, we make a summary of our results and draw some conclusions.
+The source also showed soft time lag of $0.014\pm 0.001$ sec between energy bands of $3-10$ and $0.5-3$ keV on August 29 2023 (MJD 60185) (Debnath et al.
+(2023) using NICER data).
+With reflection spectroscopy, authors found the disk inclination to be $\sim 87^{\circ}$.
+However, to date, there is no detailed study on the origin of the spectral and temporal properties of the source, which prompted us to analyze the broadband X-ray data from Insight-HXMT.
+The paper is organized as follows: in §2, we describe the observation, data selection, reduction, and analysis procedures.
+In §3, we portray the spectral and temporal properties and observed results.
+In §4, we discuss the possible physical origin of our result and connection between observed features.
+Finally, in §5, we summarize the results and draw conclusions.
 2 Observation, Data Selection, Reduction, and Analysis
-This source is being extensively monitored by various X-ray satellites, e.g., NICER, NuSTAR, Insight-HXMT, IXPE, etc.
-For our analysis, we use China’s first dedicated X-ray satellite Insight-High X-ray Modulation Telescope (HXMT; Zhang et al.
-2020).
+The source is being extensively monitored by various X-ray satellites, as discussed earlier.
+For this study, we use China’s first dedicated X-ray satellite Insight-HXMT (HXMT; Zhang et al.
+2014, 2020) data.
 In the subsequent subsections, we discuss the data selection, reduction, and analysis, respectively.
 2.1 Data Selection
-This source was regularly monitored by the Insight-HXMT satellite.
+The source was regularly monitored by the Insight-HXMT satellite.
 In the data archive, observations are available on-demand.
-At the time of the work, we found that there was a total of 34 observation IDs were available, although 3 of them were not downloadable.
-Out of those usable 31 observation IDs, we took the first 10 observation IDs.
-However, unlike other X-ray satellite data, each observation ID in HXMT has several different exposures.
-Also, some of the observation IDs continue for almost 2 to 3 days.
+We found that there are a total of 34 observation IDs available  [^1] Although 3 of them were not downloadable.
+Out of those available 31 observation IDs, we took the first 10 observation IDs.
+However, unlike other X-ray satellite data, each observation ID in HXMT has several exposures.
+Also, some of the observation IDs have continuous observations of about 2 to 3 days.
 We found that for these first 10 observation IDs, each of them has a minimum of 7 to a maximum of 18 exposures.
-We checked the analysis of all the exposures for the first 3 observation IDs and found that no significant change was noticed.
-Thus we did not fit every exposure for each observation ID.
-The details about the selected data are given in Table 1.
-Table 1: List of Data used
-Obs. Id.[1] Start Date[2] End Date[2] Start MJD[3] Stop MJD[3] Exp. (s)[4]
+Initially, we perform the analysis of all the exposures for the first 3 observation IDs and found no significant change in the parameters of spectral fits in the consecutive exposures.
+Therefore, we avoid fitting every exposure of each observation ID.
+The details of the selected data are given in Table 1.
+Table 1: List of observations used for this work.
+Here, column 1 lists the observation IDs.
+Columns 2 and 3 denote the start and end date of observations, while columns 4 and 5 refer to corresponding MJDs.
+Column 6 gives the exposure time of each observation.
+Obs. Id. Start Date End Date Start MJD Stop MJD Exp. (s)
 (1) (2) (3) (4) (5) (6)
@@
 P0614338011 2023-09-10 2023-09-11 60197.0753 60198.0667 85650
-[1] Column 1 lists each observation ID.
-[2] In columns 2 and 3, we give the start and end date of each observation ID.
-[3] In columns 4 and 5, we give the start and end MJD of each observation ID.
-[4] Column 6 gives the exposure time of each observation ID.
 2.2 Data Reduction
-Insight-HXMT, China’s first astronomical X-ray satellite, was launched on 2017 June 15 (Zhang et al.
-2014, 2020) to observe and study the properties of black holes, neutron stars (NSs), active galactic nuclei (AGNs), and other astronomical phenomena.
-The main scientific instrument is an array of 18 NaI/CSI phoswich scintillation detectors, each of which covers an effective area of 286 cm2.
+The main scientific instrument of Insight-HXMT is an array of 18 NaI/CSI phoswich scintillation detectors, each of which covers an effective area of 286 cm2.
 The satellite has three instruments or payloads: High Energy (HE; Liu et al.
@@
 After downloading on-demand level-1 data from the archive, we produced cleaned level-2 data for science analysis.
-This process of cleaning raw data was done in the following way.
+The processing of cleaning raw data was done in the following way.
 We first installed the HXMTDAS [^2] http://hxmt.org/index.php/usersp/dataan v2.05.
 Using this software, we run the hpipeline command using proper input and output directories.
-This pipeline performs several subsequent automatic commands for all three instruments.
-However, some suitable conditions are needed to be provided.
-Some conditions were set to get a good time interval, e.g., elevation angle $>10^{\circ}$; geomagnetic cutoff rigidity $>8~{}GeV$; pointing offset angle $<0.04^{\circ}$; $>600~{}s$ away from the South Atlantic Anomaly (SAA).
-All these commands altogether extract and clean the raw data and produce cleaned and analyzable science products.
-This is discussed in the HXMT manual [^2] http://hxmten.ihep.ac.cn/SoftDoc/501.jhtml in full details.
-The specific commands hespecgen, mespecgen, and lespecgen produce the spectra for HE.
-ME, and LE instruments.
-Whereas, the helcgen, melcgen, and lelcgen commands produce the light curve files for the three instruments.
-herspgen, merspgen, and lerspgen produce suitable response files.
+This pipeline performs several subsequent automatic commands for all three instruments under some given conditions.
+Some conditions were set to get a good time interval, e.g., elevation angle $>10^{\circ}$; geomagnetic cutoff rigidity $>8$ GeV; pointing offset angle $<0.04^{\circ}$; $>600$ sec away from the South Atlantic Anomaly (SAA).
+All these commands altogether extract and clean the raw data and produce cleaned and analyzable science products [^4] This is discussed in the HXMT manual [^3] http://hxmten.ihep.ac.cn/SoftDoc/501.jhtml in full details.
+The specific commands hespecgen, mespecgen, and lespecgen produce the spectra for HE, ME, and LE instruments.
+Whereas, the helcgen, melcgen, and lelcgen tasks produce the light curve files for the three instruments.
+herspgen, merspgen, and lerspgen produce corresponding response files.
 The background subtraction for both the timing and spectral data was performed using the commands hebkgmap, mebkgmap, and lebkgmap for instruments HE, ME, and LE, respectively.
-For $\chi^{2}$ fit-statistics in XSPEC, we group the spectrum using the grppha task of FTOOLS to a minimum of 30 counts per bin.
-We also set the time bin size to $0.01s$ to produce suitable light curves for generating PDS and looking for QPOs.
-Using the cleaned spectra and light curves, we analyze the source properties, which are discussed in detail in the very next subsection.
+For the $\chi^{2}$ fit-statistics in XSPEC, we grouped the spectrum using the grppha task of FTOOLS to a minimum of 30 counts per bin.
+We also set the time bin size to $0.01$ sec to produce the light curves for generating PDS and QPOs.
+Using the cleaned spectra and light curves, we further fitted and analyzed to extract the outburst properties, which are discussed in the next section.
 2.3 Data Analysis
-We perform timing and spectral analysis of the very first outburst during 2023 of the BHC Swift J1727.8-1613.
-We have downloaded the daily average light curve using the MAXI/GSC [^3] http://maxi.riken.jp/star_data/J1727-162/J1727-162.html public archive.
-Using light curves from all three modules of the HXMT, i.e., LE, ME, and HE, we have produced $0.01~{}s$ time-binned light curves first.
-Using those light curves, we have produced power density spectrum (PDS) by using the fast Fourier transformation (FFT) in the powspec task of the XRONOS package in the HEASoft software.
+We have downloaded the daily average light curve of MAXI/GSC [^5] http://maxi.riken.jp/star_data/J1727-162/J1727-162.html from the archive to study the outburst profile in addition to Insight-HXMT.
+Using the unbinned light curves from LE, ME, and HE modules, we have first produced $0.01$ sec time-binned light curves.
+Then the power density spectrum (PDS) was generated using the fast Fourier transformation (FFT) technique in the powspec task of the XRONOS package in the HEASoft software.
 Each observation’s data was subdivided into several intervals, each of which contains 8192 newbins.
 First, the PDS for each interval is generated and then they are averaged to make a resultant PDS.
-The PDS is normalized in a way that their integral will give the rms-squared fractional variability.
-White noise level is subtracted and a geometrical rebinning of -1.02 is used.
+The PDS is normalized in a way that their integral can give the RMS-squared fractional variability.
+The white noise level is subtracted and a geometrical rebinning of -1.02 is used.
 For some data, we needed to use a geometrical rebinning of -1.05 to have a prominent QPO-like nature in the PDS.
-Using these steps, we searched for LFQPOs.
-We fit the QPO shape by the use of a Lorentzian model in powspec and extract QPO properties like frequency ($\nu_{qpo}$), full-width at half maximum or FWHM, normalization.
-For many observations, we have found the presence of single or multiple harmonics.
-We have also extracted their properties by fitting them with a Lorentzian model.
-We fitted the light curves of all the exposures of the listed 10 observations.
+We fit the QPO using single or multiple Lorentzian model components in powspec to extract QPO properties like frequency ($\nu_{qpo}$), full-width at half maximum or FWHM, and normalization.
+For several observations, we have found the presence of single or multiple harmonics.
+We have also extracted their properties using the same fitting method.
+We have fitted the light curves of all the exposures of the listed 10 observations.
 We report them in the result section.
-For spectral analysis, we have also used all three modules, i.e., LE, ME, and HE and fitted the broadband data in the $2-150$ keV energy band.
-Looking for the best fit, we have found that using the combination of disk blackbody, power-law, gaussian, and pexrav models, the data is best fitted.
+For the spectral analysis, we have used the same LE, ME, and HE modules and fitted the broadband data in the $2-150$ keV energy band.
+The best is achieved by using the combination of disk blackbody, power-law, gaussian, and pexrav models.
 For some observations, we needed to add a gabs model to achieve the best fit.
 For interstellar absorption, we have used the tbabs model.
-Since we are simultaneously fitting all three modules, we have included a constant to normalize the three resultant fittings.
-The following two are our model combination: i) constant*tbabs*gabs*(diskbb + power-law + gaussian + pexrav) (for those for which gabs was required), ii) constant*tbabs*(diskbb + power-law + gaussian + pexrav) (for those for which gabs was not required.
-The spectral analysis was not done for all the exposures of the listed observations.
-Like the timing analysis, we did not spectrally analyze all the exposures, given in Table A.
-The spectrally analyzed exposures are marked with ‘*’ in Table A.
-The reason behind this is that, for timing analysis, we found variations of timing properties within a single day, whereas, for spectral analysis, the properties do not vary that much within such a short time.
+Since we simultaneously fit all three modules, we have included a constant to normalize the three resultant fittings.
+The following two model combinations are used: i) constant*tbabs*gabs*(diskbb + power-law + gaussian + pexrav), ii) constant*tbabs*(diskbb + power-law + gaussian + pexrav) (for which gabs was not required.
+The spectral analysis was done for the exposures are marked with ‘*’ in Table 2.
+We did not do the spectral analysis for all epochs as spectral variability was not present within such a small timescale (intraday) like the variability in the light curves.
 We report them in the result section.
-We note that light curves for the LE instrument were not produced for some exposure IDs although we run the same pipeline for all three instruments.
-Thus we only represented those exposures MJDs in Table 3 for which all three light curves were available for analysis for uniformity.
+We note that light curves for the LE instrument were not produced for some exposures.
+Therefore, we represented only those exposures’ MJDs in Table 3 for which all three light curves were available for analysis for uniformity.
 3 Results
 Using the timing and spectral analysis, we have studied the accretion flow properties of the very recent 2023 outburst of the BHC Swift J1727.8-1613.
-We studied the QPO properties of the source during the outburst as well as the spectral nature and radiation properties of the source using Insight-HXMT data.
-We broadly discuss the results in the two subsections below.
+We studied the QPO properties of the source during the outburst as well as the spectral nature and radiation properties using Insight-HXMT data.
+We broadly discuss the results in the next two sections.
 3.1 Temporal Properties
 Here, we discuss the evolution of the flux and hardness ratio, and QPO properties during the first 10 observation IDs of the outburst.
+3.1.1 Outburst Profile and Hardness Ratio
 Figure 1: Variation of the (a) MAXI/GSC $2-20$ keV count rate, (b) MAXI/GSC $2-6$ and $6-20$ keV count rates, and (c) hardness ratio with time.
 The HR is the ratio of the $6-20$ keV count rate to the $2-6$ keV count rate of the MAXI/GSC data.
-3.1.1 Outburst Profile and Hardness Ratio
-In Figure 1, we show the variation of the flux during the outburst.
+Figure 1 shows the variation of the flux obtained from MAXI/GSC during the outburst.
 The outburst started roughly at around MJD 60180 (2023 August 24), when its flux came out of quiescence.
@@
 Within five days, the flux increased very rapidly and the total flux reached its peak on MJD 60185 (2023 August 29).
-Then the flux started to decrease very slowly.
-It decreased very slowly and gradually except for the period from MJD 60200 to 60225, where the flux showed a significant variation.
-In panel (b), we show the variations of the $2-6$ and $6-20$ keV energy bands count rates with time.
-We notice that, after MJD 60185, the hard flux (i.e., $6-20$ keV) decreased very slowly and gradually without showing any other variation, whereas, the soft $2-6$ keV flux showed a lot of variation just like the total $2-20$ keV flux.
-MJD 60185 is the peak for both the total $2-20$ keV and hard $6-20$ keV fluxes, whereas, the soft flux reached its peak much later at around MJD 60217.
+After that, the flux started to decrease very slowly, except for the period from MJD 60200 to 60225, where the flux showed a significant change.
+In panel (b), we show the variations of the count rates with time in the $2-6$ and $6-20$ keV energy bands.
+We notice that, after MJD 60185, the hard flux (i.e., $6-20$ keV) decreased very slowly, whereas, the soft $2-6$ keV flux showed a lot of variation just like the total $2-20$ keV flux.
+On MJD 60185 the $2-20$ keV and hard $6-20$ keV fluxes reached their peak values, whereas, the soft flux reached its peak much later at around MJD 60217.
 This suggests that the variation in the total count rate is mainly due to the variation of the soft component.
-After this period, the soft flux also decreased slowly to reach the end of the outburst and enter again into the quiescence phase.
-In the panel (c), we show the variation of the hardness ratio, which is the ratio of the hard $6-20$ keV to the soft $2-6$ keV flux.
-At the start of the outburst, the HR was $\sim 0.5$.
-From this value, it first gradually decreased to MJD 60235 as the outburst progressed.
-After that, it became almost constant, although decreasing very slowly to $\sim 0.025$ at the end of our taken data of the outburst on MJD 60350.
-The hardness ratio is defined as $F_{H}/F_{S}$, where $F_{H}$ and $F_{S}$ are the hard and soft fluxes respectively.
-When this value is on the higher side, the spectral nature of the source is hard.
+After this period, the soft flux also decreased slowly till the end of the outburst and entered again into the quiescence phase.
+In panel (c), we show the variation of the hardness ratio, which is the ratio of the hard ($F_{H}$) $6-20$ keV to the soft ($F_{S}$) $2-6$ keV flux.
+At the start of the outburst, the HR was $\sim 0.5$ and then it gradually decreased to MJD 60235 as the outburst progressed.
+After that, the HR became almost constant, although decreasing very slowly to $\sim 0.025$ at the end of the considered data of the outburst on MJD 60350.
+The way HR is defined, when this value is on the higher side, the spectral nature of the source is hard.
 This is the general case when an outburst is starting.
 When this value is low, there is dominance of soft flux over hard flux and the spectral nature would be soft.
-Without studying the detailed spectral analysis, HR provides a quick and rough idea about the spectral nature of an outburst.
+Without studying the detailed spectral analysis, HR provides a quick idea about the spectral nature of an outburst.
 However, for this source, we notice that HR was a little higher at the onset of the outburst, where the hard flux was also comparable to the soft flux.
@@
 Since we are not analyzing the full duration of the outburst, it is not the scope of the paper to study the full spectral classification.
-We only spectrally define the source in the duration which falls into our analysis range.
+We only define the spectral state of the source based on the duration of which data is used in this work.
 3.1.2 Quasi Periodic Oscillations (QPOs)
-As described in §2, we used the $0.01s$ time-binned light curves from all three bands (LE, ME, and HE) for our timing analysis.
-Using those light curves, we produced the power density spectrum (PDS) of those light curves to study QPOs.
-We find that QPO nature was present in all of the light curves in all three energy bands.
+As described in §2, we used the $0.01$ sec time-binned light curves from all three bands (LE, ME, and HE) and produced the PDS to study QPOs.
+We find that QPO nature was present in all the light curves in all three energy bands.
 In Figure 2, we show the variation of the QPO frequency ($\nu_{qpo}$) for the 10 different obs.
-IDs that we have used for our timing analysis.
+IDs as stated before.
 We first checked the first obs.
@@
 It has several exposure IDs (given in Table 2).
-While we fit every exposure ID of this obs ID, we found that the QPO frequency was changing even within this small period.
-This is why we fitted all the light curve data of the exposures of the obs IDs that we have taken for analysis.
+While we fit every exposure of this obs ID, we found that the QPO frequency was changing even within this small period.
+This is the reason we fitted all the light curves of the exposures of the obs IDs that we have taken for analysis.
 We find that for every single obs ID, the $\nu_{qpo}$ showed significant variation.
@@
 Evolution of QPO frequency in different observation ID.s
-Refer to captionRefer to caption
+Refer to captionRefer to captionRefer to caption
 Figure 2: Evolution of QPO frequencies (Hz) with time.
@@
 After this date, it increased very slowly till MJD $\sim 60190$ when $\nu_{qpo}$ was $\sim 1.45$.
-Then it decreased for a very short period until it very rapidly increased to $1.5$Hz again in all three bands.
-After that, it showed varying nature till MJD $\sim 60197$ and then it finally increased again.
-This might be caused by the varying nature of the origin of these oscillations.
-We will discuss this in the next section.
+Then it decreased for a very short period until it rapidly increased to $1.5$Hz again in all three bands.
+After that, it showed varying nature till MJD $\sim 60197$, and then it finally increased again.
+We have discussed the possible reason behind this nature in the next section.
 Using the PDS fitting, we extracted some information (e.g., full-width at half maximum (FWHM), Normalization (LN)) of QPOs.
-We also extracted source and background count rates from the exposures.
+We also extracted source and background count rates for the exposures.
 Using these estimates, we calculated the $Q$-value and RMS that represent the sharpness of the QPO and the fractional variability in the PDS, respectively.
-The values are listed in Table 3 in columns 5–7 ($Q$-value) and 8–10 (RMS) for LE, ME, and HE respectively.
+These values are listed in Table 3 in columns 5–7 ($Q$-value) and 8–10 (RMS) for LE, ME, and HE respectively.
 Evolution of RMS (%) with QPO frequency and time
@@
 Figure 4: 3-D representation of the variation of $\nu_{qpo}$ and (%)RMS with time (MJD) for (a) LE, (b) ME, and (c) HE.
-In Figure 4, we show a 3-D representation of the evolution of QPO RMS with corresponding frequency.
-In the LE band (Fig.4a), The RMS was higher at the start of the analysis period and it then gradually decreased.
+In Figure 4, we show a 3-D representation of the evolution of QPO and RMS with time.
+In the LE band (Fig.
+4a), the RMS was higher at the start of the analysis period and it then gradually decreased.
 In the ME band (Fig.
 4b), it increased from the onset of the analysis period for some time and then again decreased.
-HE band (Fig.
-4c) shows a similar trend as of LE band.
+The HE band (Fig.
+4c) shows a similar trend as in LE band.
 Since the QPO nature has already been reported in the recent paper by Yu et al.
-(2024), we would not put much insight into the classification of the QPO nature.
-However, by the values of both $Q$-factor and RMS, it can be said that the QPO is type-C in nature for all the exposure IDs that have been presented here.
+(2024), we are not focusing on the classification of the QPO nature.
+However, by the values of both $Q$-factor and RMS, it can be said that the QPO is type-C in nature for all the exposures reported here.
 3.2 Spectral Properties
-Along with studying the timing properties, studying the spectral properties gives a clearer idea about the nature of the source.
-With the available Insight-HXMT data, we analyzed the source for a total of 32 exposure IDs.
+Along with the timing properties, studying the spectral properties provides more insights on the nature of the outburst.
+With the available Insight-HXMT data, we analyzed the source for a total of 32 exposures.
 These exposure IDs are marked with a ‘*’ sign in the 1st column of Table 2.
-This source hasn’t been extensively analyzed spectrally.
-We perform an extensive study of this source’s spectral properties for the first few days after the onset of its outburst.
-We started our spectral analysis on MJD 60181.4.
-We have used LE+ME+HE in the $2-150$ keV energy band (LE in 2–10, ME in 10–35, and HE in 35–150 keV) for our spectral fitting for all these exposure IDs.
+An extensive spectral analysis on this source has not been done so far.
+We perform a detailed spectral analysis on this source for the first time using HXMT data.
+We started our spectral analysis from MJD 60181.4.
+We have simultaneously used LE+ME+HE in the $2-150$ keV energy band (LE in 2–10, ME in 10–35, and HE in 35–150 keV) for our spectral fitting for all the selected exposure IDs.
 Model Fitted Insight-HXMT Spectra
@@
 (a) is or observation ID.
-P061433800501 (Exposure: P061433800501-20230902-01-01) for which the gabs model was necessary to achieve the best fit.
+P061433800501 (Exposure: P061433800501-20230902-01-01) for which the gabs model was required to achieve the best fit.
 (b) is for the observation ID.
-P061433800601 (exposure: P061433800601-20230904-01-01) for which the gabs model was not needed to achieve the best fit.
+P061433800601 (exposure: P061433800601-20230904-01-01) for which the gabs model was not required to achieve the best fit.
 We first fitted the data with phenomenological diskbb and power-law models.
@@
 Besides these features, the absorption feature was noticed below 2 keV.
-Thus, we refitted using the model combination as constant*tbabs*gabs(diskbb + power-law + ga + pexrav).
+Thus, we refitted the data using the model combination as constant*tbabs*gabs(diskbb + power-law + ga + pexrav).
 Here, the gabs model takes care of the low energy absorption feature, while the ga and pexrav models take care of the contribution from the Gaussian and hump-like nature, as mentioned above.
@@
 P061433800601 (exposure ID: P061433800601-20230904-01-01), for which $\chi^{2}/DOF\sim 0.9$.
-Model Fitted Insight-HXMT Spectra
+Model Fitted Parameters
 Refer to captionRefer to caption
@@
 Here, $\Gamma$ is the spectral fitted index using combined LE+ME+HE data.
-The QPO frequency could be produced from the oscillation of the Compton corona, which could also be the origin of the hard power-law photons that extend the spectrum to very high energies.
-They could be correlated to each other.
-It has been previously seen that the QPO frequency ($\nu_{qpo}$) and the photon index of power-law ($\Gamma$) show a positive correlation for some of the sources during their active outbursting phase (Vignarca et al.
+The QPO frequency can be produced from the oscillation of the Compton corona, which can also be the origin of the hard power-law photons that extend the spectrum to higher energies.
+Therefore, these two features might be interlinked or correlated.
+It has been previously seen that the QPO frequency ($\nu_{qpo}$) and $\Gamma$ show a positive correlation for some of the sources during their active outbursting phase (Vignarca et al.
 2003; Shaposhnikov & Titarchuk 2009; Stiele et al.
@@
 Considering these points, we have found a positive correlation between $\Gamma$ and $\nu_{qpo}$.
-Using both the Pearson-Linear and Spearman-Rank correlation methods, we found correlation coefficients for the three energy bands as below.
+Using both the Pearson-Linear and Spearman-Rank correlation methods, we found correlation coefficients for the three energy bands as shown in the table below.
 Pearson-Linear Spearman-Rank
@@
 4 Discussions
-The Galactic black hole Swift J1727.8-1613 was monitored daily by various X-ray satellites like NICER, NuSTAR, Insight-HXMT, etc.
-Multi-wavelength follow-up observations were also carried out for this source.
-We have taken HXMT data for both our spectral and timing analysis from MJD 60181.4 to 60198.0.
+The Galactic black hole Swift J1727.8-1613 went through an outburst recently in 2023.
+We have used Insight-HXMT data for our both spectral and timing studies from MJD 60181.4 to 60198.0.
 The evolution of the light curve makes the outburst nature of the source a fast-rise slow decay type.
-After the onset of the outburst, it rose to its peak very fast within 5 days.
-Using the $0.01~{}s$ time-binned light curves from the three instruments of HXMT (LE, ME, and HE), we studied the source’s timing properties by analyzing the light curves.
-Using the spectra files from these three instruments we also studied the spectral properties of this source by analyzing combined LE+ME+HE spectra in the $2-150$ keV broad energy range.
+After the onset of the outburst, the flux reached its peak value very fast within 5 days.
+Using the $0.01$ sec time-binned light curves from the three instruments of HXMT (LE, ME, and HE), we studied the source’s timing properties.
+Using the spectra files from these three instruments, we also studied the spectral properties of this source by analyzing combined LE+ME+HE spectra in the $2-150$ keV broad energy range.
 4.1 Propagation of Shock
@@
 Quasi-periodic oscillation is one of the most important and common phenomena for stellar mass black holes.
-For this newly discovered source, we analyzed light curves of a total of 342 light curves (114 for LE, ME, and HE each).
+For this newly discovered source, we analyzed a total of 342 light curves (114 for LE, ME, and HE each).
 However, in some of the exposures, light curves were not produced properly.
@@
 We list those data in Table 3.
-Our analysis is final for 276 data.
-We found the presence of QPO in all of these light curves and they show rapid evolution in frequency in this short analysis period.
+Finally, we analyzed 276 lightcurves.
+We found the presence of QPO in all of these light curves, and they showed rapid evolution in frequency in this short analysis period.
 The QPO frequency ($\nu_{qpo}$) showed an increase in frequency within even a single day.
-This is one of the advantages of using Insight-HXMT as it provides multiple exposures for one observation ID which helps to study the short-scale variations of such important physical properties of the source.
 The evolution of the QPO frequency has been discussed broadly in the result section.
@@
 From the variation of QPO frequency, (%)RMS, and $Q$-factor, the nature of the QPO is designated to be type-C.
-Although this is a well-established topic in the literature, the origin of QPO is still a matter of debate.
-Many models try to explain the type-C QPO phenomenon, e.g., the Lense-Thirring effect (Stella & Vietri 1998), accretion-ejection instability (Tagger & Pellat 1999), oscillation of the disk (Titarchuk & Osherovich 2000), the misalignment between the spin and binary rotational axis (Ingram et al.
-2009, and references therein).
-We would like to focus on a physical scenario that the QPOs could have originated from the shock instabilities in advective flows around black holes.
-Black hole accretion is transonic in which multiple sonic points could exist (Chakrabarti & Titarchuk 1995).
-Depending on the flow properties, there is the possibility of the formation of standing shocks (Singh et al.
-2022 and references therein).
+Although the QPOs are extensively studied in the literature, their origin requires in depth modeling.
+Here, we would like to focus on the physical scenario that explains the origin of QPOs due to the shock instabilities in advective flows around BHs.
+Accretion onto BH is transonic flow with possibility of multiple sonic points (Chakrabarti 1989).
 The matter from the companion need not necessarily be only Keplerian.
@@
 Such a lower angular momentum component of infalling matter gets accreted in a free-fall timescale.
-Due to the tug of war between the pressure of gravitational force and centrifugal force, this matter could virtually stop from falling onto the BH at some distance from the BH and undergo shock transition, creating a post-shock region.
-This approach could very well describe the spectral and temporal properties of BH, as it has been shown previously in both theoretical (Chakrabarti & Titarchuk 1995) and observational (Debnath et al.
+Due to the tug of war between the gravitational force and centrifugal force, this matter could virtually stop at some distance from the BH and undergo shock transition, creating a post-shock region.
+Depending on the flow properties, the standing shocks forms (Chakrabarti 1989; Singh et al.
+2022 and references therein).
+This approach can well describe the observed spectral and temporal properties of BHs (Debnath et al.
 2014; Mondal et al.
 2014, Chatterjee et al.
-2020, 2021, 2023) studies.
-The change in size of such post-shock region can describe the spectro-temporal properties during an outburst.
+2020, 2021, 2023).
 According to Molteni et al.
-(1996), the oscillation in the shock region could have originated due to the presence of cooling.
-This was later shown in simulation works by Garain et al.
+(1996), the shock can be oscillatory due to the presence of cooling.
+This was later shown in simulation by Garain et al.
 (2014).
-The oscillation of the shock region could have originated due to the resonance condition between the cooling timescale due to the Comptonization process and the infall timescale in the coronal region (see Chakrabarti et al.
+During the oscillation when the cooling timescale due to the Comptonization process and the heating timescale matches, QPOs originate (see Chakrabarti et al.
 2015).
 Additionally, it has also been observed by the authors that once the QPO sets in, it gets locked for a week or more depending on the above condition.
-From the expressions of infall and cooling timescales, we get that the QPO frequency ($\nu_{qpo}$) could be given as (Molteni et al.
+From the expressions of infall and cooling timescales, we get that the QPO frequency ($\nu_{qpo}$) can be estimated using the relation (Molteni et al.
 1996; Chakrabarti & Manickam 2000; Chakrabarti et al.
 2004),
-$\nu_{qpo}=\frac{c^{3}}{2GM_{BH}}\frac{1}{RX_{s}(X_{s}-1)^{1/2}}Hz,$ (1)
-where $c$ and $G$ are the speed of light and gravitational constant respectively.
-$M_{BH}$ represents the mass of the black hole, $X_{s}$ is the distance of the shock, or called shock location.
-$X_{s}$ is in the units of Schwarzschild radius ($r_{s}=2GM_{BH}/c^{2}$).
-$R$ is the ratio of matter densities in post-shock to pre-shock regions ($\rho_{+}/\rho_{-}$) and is known as the compression ratio.
+Where $c$, $G$, $M_{BH}$, $X_{s}$, and $R$ are the speed of light, gravitational constant, the mass of the BH, location of the shock, and the ratio of matter densities in post-shock to pre-shock regions ($\rho_{+}/\rho_{-}$) respectively.
 From our timing analysis, we already extracted the information about QPO frequency ($\nu_{qpo}$).
@@
 According to Chakrabarti et al.
-(2005), $X_{s}$ could be located anywhere from 10 to 1000 $r_{s}$.
-The shock stays very far away $\sim 1000~{}r_{s}$ when the spectral nature of the ongoing outburst is hard and becomes smaller gradually in the process of cooling.
-Using the relation (1), we found that the shock was far at the start of the analysis period (Fig.
-8b).
-It then rapidly decreased as the QPO frequency increased rapidly during the initial few days.
-Close to MJD 60185, when the $\nu_{qpo}$ stopped its rapid increment, and the shock location started to decrease very slowly.
+(2005), $X_{s}$ can be located anywhere above 10 $r_{s}$ depending on the flow parameters producing shocks.
+The shock forms far away $\sim 1000~{}r_{s}$ when the spectral nature of the ongoing outburst is hard and decreases gradually in the progressive days as cooling increases (Mondal et al.
+2015).
+Using Eq.
+1, we found that the shock was far away from the BH at the start of the outburst (see Figure 8b).
+It then significantly decreased as the QPO frequency increased rapidly during the initial few days.
+Close to MJD 60185, when the $\nu_{qpo}$ stopped increasing, and the shock location started to decrease very slowly.
 The values of the shock location along with QPO frequency are given in Table 3.
-As equation (1) suggests, as the shock was bigger at the start of the outburst, the QPO frequency was small.
-As it decreased rapidly due to cooling, the $\nu_{qpo}$ increased rapidly.
-Then when the cooling was less efficient, the size of the shock changed very slowly, although decreasing.
-It made the $\nu_{qpo}$ increase with time but not so rapidly as before.
 If the time of the first exposure ID was taken to be 0, and the shock location as $X_{s0}$ and after time $t$, if the shock is at $X_{s}$, then the velocity of the movement of the shock would be given by the equation (Chakrabarti et al.
@@
 […]
-Using this relation, we estimated the velocity of shock.
-This is given in Fig.
-8c.
+Using this relation, we have estimated the velocity of the movement of the shock.
+The estimated shock velocities are given in Figure 8c, which is a few tens of m/s in accord with the previous estimations (Chakrabarti et al.
+2005; Mondal et al.
+2015).
 We find that the shock velocity has varied within this period.
-In the beginning, it attained higher values when the $X_{s}$ was also larger, however decreasing rapidly.
+At the beginning, it attained higher values when the $X_{s}$ was also larger, however decreasing significantly afterward.
 It slowed down as the shock moved inwards, along with some fluctuations.
-This could be due to the limitation of assuming constant velocity of the movement of the shock.
+This can be due to the limitation of assuming constant velocity of the shock propagation.
 4.2 Evolution of the physical configuration and spectral nature
-From the spectral properties, we have seen that the $T_{in}$ was first low and it gradually increased.
+From the spectral properties, we have seen that the $T_{in}$ was low initially and then gradually increased.
 The exact opposite trend was observed for the Normalization of the diskbb model.
@@
 $i$ is the inclination angle of the accretion disk to the observer.
-Considering, $D_{10}$ and $\theta$ do not change during an outburst, the relation could be shortened like $R_{in}\sim N_{diskbb}^{1/2}$.
-Thus as the $N_{diskbb}$ decreased, it means the inner edge of the disk moved closer to the black hole.
-This is also supported by the fact that $T_{in}$ increased as the disk came closer.
-This also supports the properties we showed from the timing analysis.
-The shock forms at a distance ($X_{s}$) when the sub-Keplerian matter is halted due to the counterbalance between the pressures of gravitational force and the centrifugal force.
-This forms the hot Compton cloud, as it becomes hot and puffed up in time.
-The hot Compton cloud will intercept some of those disk photons and there will be scattering in the process of inverse-Comptonization that will increase the energy of those intercepted photons and will emit them at higher energies.
-In this way, the Compton cloud will be cooled due to energy loss.
-Due to this cooling, the Compton cloud will gradually shrink in size.
-As it continues shrinking in size, the disk will continue to move inwards.
-As we have seen from the timing properties, the shock location was larger at the start of the outburst.
-The inner-disk radius was also higher as $R_{in}\sim N_{diskbb}^{1/2}$ and $N_{diskbb}$ was higher.
-Then cooling starts as the disk starts to approach inwards.
-In this way, the Compton cloud was losing energy by up-scattering and as the cooling continued, the $X_{s}$ and $R_{in}$ were both moving closer to the black hole.
+Assuming that $D_{10}$ and $\theta$ do not change during an outburst, the relation could be shortened like $R_{in}\sim N_{diskbb}^{1/2}$.
+For our case, except for one exposure, the inclination varied in a narrow range of 78-86.
+Thus, as the $N_{diskbb}$ decreased, it means the inner edge of the disk moved closer to the BH, which agrees with the fact that $T_{in}$ increased as the disk moved closer.
+Furthermore, this also supports the properties we showed from the timing analysis that the shrinking of post-shock region with time.
 Looking at the spectral parameters, and also the location of the shock, we designate that the source was already in an intermediate state at the start of our analysis period.
 As the outburst progressed, it was transitioning towards softer states.
-From the values of the parameter $\cos(i)$ of the pexrav model, we find that the inclination of the disk was $>60^{\circ}$, which makes this source a high inclination binary system.
+From the values of the parameter $\cos(i)$ of the pexrav model, we find that the inclination of the disk was $>60^{\circ}$, which makes this source a highly inclined system.
 The hydrogen column density ($N_{H}$) showed a variation, which is common for Galactic black hole binaries unless the system is heavily obscured, which has recently been found for the newly discovered BH Swift J151857.0-572147 (Mondal et al.
@@
 (v) The spectral and temporal properties correlate to each other which makes our claim stronger about the origin of QPOs.
-(vi) This source is in a high inclination binary system.
+(vi) This source has a high inclination.
 (vii) The hydrogen column density varied in the range of $0.10-0.35\times 10^{22}$ cm-2, which is quite common for interstellar absorption of Galactic black holes.

### v3

@@
 Abstract
-The transient Galactic black hole candidate Swift J1727.8-1613 went through an outburst for the very first time that started in August 2023 and lasted for almost 6 months.
+The transient Galactic black hole candidate Swift J1727.8-1613 went through an outburst for the very first time in August 2023 and lasted for almost 6 months.
 We study the timing and spectral properties of this source using publicly available archival Insight-HXMT data for the first 10 observation IDs that last from MJD 60181 to 60198 with a total of 92 exposures for all three energy bands.
-We extracted the quasi-periodic oscillation properties by model fitting the power density spectrum and from those properties we designate that the QPOs are type-C in nature.
+We have detected quasi-periodic oscillations (QPOs) in a frequency range of $0.21\pm 0.01$ $1.86\pm 0.01$ Hz by fitting the power density spectrum.
+Based on the model-fitted parameters and properties of the QPOs, we classify them as type-C in nature.
 We also conclude that the origin of the QPOs could be the shock instabilities in the transonic advective accretion flows around black holes.
@@
 interstellar absorption, power-law, multi-color disk-blackbody continuum, Gaussian emission/absorption, and reflection by neutral material.
-From the spectral properties, we found that the source was in an intermediate state at the start of the analysis period and was making a transition toward the softer states.
+From the spectral properties, we found that the source was in an intermediate state at the start of the analysis period and was transitioning to the softer states.
 The inner edge of the accretion disk moved inward in progressive days following the spectral nature.
-We found that the source has a high inclination.
-The average hydrogen column density estimated from the model fitting is $0.27_{-0.17}^{+0.08}\times 10^{22}$ cm-2.
-X-rays: binaries – stars: black holes – stars: individual (Swift J1727.8-1613) – accretion, accretion disk – shock – radiation
-journal: AAS
+We found that the source has a high inclination of $78^{\circ}-86^{\circ}$.
+The hydrogen column density from the model fitting varied in the range of $(0.12\pm 0.02-0.39\pm 0.08)\times 10^{22}$ cm-2.
+X-rays: binary stars (1811); black holes (162); Stellar accretion disks (1579); Shocks (2086); Compact radiation sources (289)
 1 Introduction
@@
 Depending on the mass of the companion star, they can be mainly of two types: low-mass X-ray binaries (LMXRBs) or high-mass X-ray binaries (HMXRBs).
-The LMXRBs consist of type-A or later type of stars, whereas the HMXRBs consist of giant O or B-type of stars (White et al.
+The LMXRBs consist of type-A or later type of stars, whereas the HMXRBs consist of giant O or B-type of stars (e.g., White et al.
 1995; Tetarenko et al.
 2016).
-During the transfer of mass from the companion, matter accumulates at a distance from the BH due to insufficient viscosity to drive the matter inward, which is called the pile-up radius (Chakrabarti et al.
+During the transfer of mass from the companion, matter accumulates at a distance from the BH due to insufficient viscosity to drive the matter inward (Mineshige 1993; Kuulkers et al.
+1997; Parmar et al.
+1997; Lasota 2001), which is called the pile-up radius (Chakrabarti et al.
 2019; Chatterjee et al.
 2022).
-Due to the accumulation of more matter, the temperature will rise.
+Due to the accumulation of more matter, the temperature of the disk rises.
 The gradual increase in temperature develops instability at the pile-up radius, which further increases the viscosity.
@@
 Thus, these systems make ideal laboratories for studying the physics of accretion.
-The BHXRBs mostly stay in a quiet dormant state when their flux remains at a very low detection level.
-Episodically, the radiation level increases by a very large factor and they become easily detectable, which is known as an outburst (Remillard & McClintock 2006).
+BHXRBs mostly stay in a quiet dormant state when their flux remains at a very low detection level.
+Episodically, the radiation level increases by a very large factor and they become easily detectable, which is known as an outburst (Tanaka & Shibazaki 1996; Remillard & McClintock 2006).
 The X-ray luminosity increases by several orders of magnitude during an outburst when compared to a quiescence.
@@
 They stay in the dormant state for most of their lifetimes.
-The luminosity becomes $\sim 10^{35}$ erg s-1 during episodic outbursting phase.
+The luminosity becomes $\sim 10^{37-38}$ erg s-1 during episodic outbursting phase (Tanaka & Shibazaki 1996).
 During the onset of an outburst, the flux in the light curve changes noticeably.
@@
 When a BH outburst goes through all these states, it is known as a type-I outburst.
-When the soft state is absent, it is known as a ‘failed’ or type-II outburst.
-The radiation spectrum of a BHXRB generally consists of two main components.
+When the soft state is absent, it is known as a ‘failed’ or type-II outburst (Hynes et al.
+2000; Brocksopp et al.
+2001; Belloni et al.
+2002; Brocksopp, Bandyopadhyay & Fender 2004; Capitanio et al.
+2009; Curran & Chaty 2013; Del Santo et al.
+2016; Tetarenko et al.
+2016; Garcia et al.
+2019; Alabarta et al.
+2021).
+The energy spectrum of a BHXRB generally consists of two main components.
 One is the soft thermal multi-colour blackbody component, which is due to the blackbody radiation of the seed photons in the standard accretion disk (Shakura & Sunyaev 1973; Novikov & Thorne 1973).
 The other is the hard power-law component, which can extend up to very high energies.
-The origin of this component is due to the inverse Comptonization of a fraction of seed photons from the disk, which is intercepted by a hot Compton cloud (Sunyaev & Titarchuk 1980, 1985; Haardt & Maraschi 1993; Zdziarski et al.
+The origin of this component is thought to be due to the inverse Comptonization of a fraction of seed photons from the disk, which is intercepted by a hot Compton cloud (Sunyaev & Titarchuk 1980, 1985; Haardt & Maraschi 1993; Zdziarski et al.
 1993; Titarchuk 1994; Chakrabarti & Titarchuk 1995; Życki et al.
 1999).
-When the outburst is in HS, there is a very small contribution from the thermal blackbody component and mostly the (non)thermal powerlaw component dominates.
-In SS, the thermal blackbody component becomes dominant while in the intermediate states (HIMS and SIMS), the contribution from the two components stays comparable.
+When the outburst is in the HS, there is a very small contribution from the thermal blackbody component and mostly the (non)thermal powerlaw component dominates.
+In the SS, the thermal blackbody component becomes dominant while in the intermediate states (HIMS and SIMS), the contribution from the two components stays comparable.
 Other than these two components, in the case of a high soft state (HSS), there could be the presence of bulk motion Comptonization (BMC) which takes place due to the relativistic speed of matter when it reaches a closer to the BH (Blandford & Payne 1981a,b; Payne & Blandford 1981; Chakrabarti & Titarchuk 1995; Psaltis & Lamb 1997; Borozdin et al.
@@
 As a result, the Fe $K\alpha$ emission line at $\sim 6.4$ keV and a reflection hump above 20 keV may originate.
-Therefore, the presence of complex features in the spectrum in BHXRBs makes them important to study to understand their origin.
+The presence of these complex features in the spectrum of BHXRBs makes it interesting to study.
+It is important to study and model these features to understand their origin.
 This can be achieved by fitting broadband data using multicomponent models.
@@
 2020) can be very useful.
-Besides rich spectral features and variabilities, temporal properties are equally important to understanding the dynamics of the accreting gas around the BHs.
-During an outburst, it has been observed that the light curves show very small timescale variabilities, especially in the high-energy bands.
-The Fourier transformation of the light curve imprints such variabilities as both broadband noise and narrow features in the power density spectrum or PDS (van der Klis 1989).
-The broadband noise is spread over a large frequency range modeled by power-law function.
-The narrow feature is a power peak in narrow frequency ranges, known as the quasi-periodic oscillation (QPO), modeled by one or multiple Lorentzian profiles.
-In BHXRBs, the low-frequency QPOs (LFQPOs) are often observed mainly due to their geometrical origin.
-Depending on their properties like the frequency ($\nu$), $Q$-value ($=\nu/\delta\nu$, where $\delta\nu$ is the full width at half maximum or FWHM), (%)RMS, etc., LFQPOs are classified into three different types: type A, B, and C (Casella et al.
+Besides rich spectral features, temporal properties are equally important to understanding the dynamics of the accreting gas around the BHs.
+During an outburst, it has been observed that the light curves show changes in very small timescale, especially in the high-energy bands.
+The Fourier transformation of the light curve imprints such changes as both broadband noise and narrow features in the power density spectrum or PDS (van der Klis 1989).
+The broadband noise is spread over a large frequency range modeled by power-law or broad Lorentzian functions.
+The peak-like feature is a power peak in narrow frequency ranges, known as the quasi-periodic oscillation (QPO).
+The PDS is modeled by one or multiple Lorentzian profiles along with the power-law model.
+The Lorentzian model helps estimate the properties of the QPO (Nowak 2000; Belloni et al.
+2002).
+Low frequency quasi periodic oscillations (LFQPOs) lie in the frequency range of a few mHz to $\sim 30$ Hz (Belloni et al.
+2002; Casella et al.
+2005; Motta et al.
+2015).
+Depending on their properties, e.g., the frequency ($\nu$), $Q$-value ($=\nu/\delta\nu$, where $\delta\nu$ is the full width at half maximum or FWHM), (%)RMS, etc., LFQPOs are classified into three different types: type A, B, and C (Casella et al.
 2005).
-Although their origin is still a topic of debate, several ideas were put forward e.g., Lense-Thirring precession (Stella et al.
+Several ideas were put forward to explain the LFQPO phenomenon in BHs.
+Although type-C LFQPO is now a well-studied phenomenon, there is still debate about its origin, as several models can explain it.
+Some of the often used explanations are the Lense-Thirring precession (Stella et al.
 1999; Ingram et al.
 2009), magneto-acoustic waves (Titarchuk et al.
-1998), accretion-ejection instability (Tagger & Pellat 1999), shock oscillation model (Molteni et al.
+1998), accretion-ejection instability (Tagger & Pellat 1999) and the shock oscillation model (Molteni et al.
 1996; Chakrabarti et al.
 2005, 2008, 2015).
-The shock oscillation model can explain both the spectral and temporal properties simultaneously, which makes it a more comprehensive model.
-As both the properties are coming from the same system, and changes in spectral states can also dictate the types of QPOs, it is believed that both are interlinked.
-A good correlation is observed between the two using just the light curve properties, e.g., the hardness ratio or HR, hardness intensity diagram, or HID (Homan et al.
-2001).
-Some model-fitted approaches also showed that the changes in both properties are due to the interplay between different mass accretion rates or variation in viscosity, can be understood through accretion rate ratio intensity diagram or ARRID (Mondal et al.
+However, compared to type-C, the other two types’ origin is still not clearly known.
+The shock oscillation is one of the features of the two-component advective flow (TCAF; Chakrabarti & Titarchuk 1995) model.
+This model suggests that there are two different forms of angular momentum distribution in matter supplied from the companion star, Keplerian and sub-Keplerian.
+Due to its high viscosity, the Keplerian component moves in a viscous timescale and forms a thin accretion disk on the equatorial plane.
+The accretion disk moves inwards as this matter gains critical viscosity.
+Less viscous accretion flow characterizes the sub-Keplerian one, which falls radially in a free fall timescale.
+This matter resides above and below the Keplerian disk.
+The sub-Keplerian component, which is optically thin, creates a shock front where infalling matter virtually stops because of the tug of war between centrifugal and gravitational forces.
+The boundary layer, sometimes referred to as the CENBOL (centrifugal pressure-supported boundary layer, Chakrabarti & Titarchuk 1995), is the shock front.
+In this model, the post-shock region is the ‘Compton’ cloud region, the repository of hot electrons.
+As blackbody photons are produced by the disk, the soft multicolor blackbody component can be explained by the Keplerian component.
+Some of the captured soft disk photons are upscattered and released as hard power-law photons by the CENBOL, which serves as the Comptonizing region.
+During the start of an outburst, the CENBOL forms farther away from the BH, and the disk is also truncated at a very large radius.
+As a result, photons originating from the disk become hard through inverse Compton scattering by hot electrons, and thus we always observe a hard state during the onset of an outburst from BHs.
+As time progresses, the Keplerian disk moves inwards, producing more blackbody radiation.
+More soft photons are intercepted in the CENBOL, which are upscattered and help to cool down the CENBOL.
+More photons are emitted in the whole process and flux increases and the source goes toward a softer state.
+In a high soft state, the CENBOL is totally quenched, and the disk reaches very close to the black hole.
+Thus, only the disk photons now contribute to the radiation, which makes the spectrum soft.
+Besides describing the spectral features and their changes, this model also explains the QPO properties.
+In the HS, when the CENBOL is bigger, it could produce LFQPOs due to the resonance oscillation of the shock, when the infall and cooling timescales match (Molteni et al.
+1996, Chakrabarti et al.
+2015, Garain et al.
+2014).
+As the outburst progresses, CENBOL loses its size, and we see a rise in the frequency of the QPO.
+This is discussed with mathematical relation in the discussion section.
+As the SS approaches, the CENBOL shrinks in size, and there is no oscillation, cooling takes over the heating timescale.
+Thus, we do not see any QPO in the SS.
+In between the HS and SS, when the CENBOL is intermediate in size, there could be the presence of either type-C or type-A/B QPOs.
+In the HIMS, we see type-C, whereas, in the SIMS, mostly type-A/B is seen.
+As changes in both the spectral and timing properties come from the same system and changes in spectral states can also dictate the types of QPOs, there could be a possible correlation between the spectral and temporal properties.
+A connection between these two properties is observed using a hardness intensity diagram, or HID (Homan et al.
+2001), which shows the variation of the flux in the light curve (considered here as timing property) with the hardness ratio, or HR, which shows roughly the variation of the spectral nature of the outburst (considered as spectral property here).
+A detailed outburst study showed that the changes in both properties are due to the interplay between different mass accretion rates or variation in viscosity, which can be understood through the accretion rate ratio intensity diagram or ARRID (Mondal et al.
 2014; Jana et al.
@@
 2020).
-The RMS-intensity diagram or RID (Munoz-Darias et al.
+The RMS-intensity diagram (RID, Munoz-Darias et al.
 2011), and hardness ratio-intensity diagram or HRD (Belloni et al.
-2005) can also make a link between spectral and temporal properties from pure observational ground.
+2005) can also link spectral and temporal properties from a pure observational ground.
+As shown in Chakrabarti et al.
+(2015), when the cooling timescale is within 50% of the infall/heating timescale, the QPO originates, which is in accord with the spectral model fitted ARRID in Mondal et al.
+(2014), where QPOs disappeared for ARR $<0.5$ in the ARRID.
+Unlike most of the explanations in the literature, the shock oscillation model in TCAF paradigm is successful in addressing the spectral and temporal properties responsible for the origin of QPOs, which prompted us to consider this model for rest of the paper to explain QPOs observed in Swift J1727.8-1613 during Insight-HXMT era.
 The BHC Swift J1727.8-1613 was discovered very recently on 24 August 2023 or MJD 60180 (Kennea & Swift Team 2023).
@@
 2023).
-There is X-ray monitoring of the source by NICER, NuSTAR, Insight-HXMT, IXPE, etc.
+There have been X-ray monitoring of the source with NICER (O’Connor et al.
+2023), NuSTAR (Dovciak et al.
+2023), Insight-HXMT (Peng et al.
+2024), IXPE (Veledina et al.
+2023), AstroSat (Katoch et al.
+2023).
 Using IXPE data on September 7, 2023, Veledina et al.
@@
 (2023).
-At the starting of the outburst, the spectral nature was consistent with a hard state with a power-law photon index ($\Gamma$) $\sim 1.5-1.7$ (Liu et al.
+At the begining of the outburst, the source was consistent with being in the hard state, described by a power-law with a photon index ($\Gamma$) $\sim 1.5-1.7$ (Liu et al.
 2023).
 Later, using the NICER observations, Bollemeijer et al.
-(2023) reported the softening of the spectrum with a substantial contribution of the soft blackbody component after 25 days from the starting of the outburst.
+(2023) reported the softening of the spectrum with a substantial contribution of the soft blackbody component after 25 days from the start of the outburst.
 The presence of LFQPOs was reported by Draghis et al.
@@
 (2023) using NICER data).
-With reflection spectroscopy, authors found the disk inclination to be $\sim 87^{\circ}$.
-However, to date, there is no detailed study on the origin of the spectral and temporal properties of the source, which prompted us to analyze the broadband X-ray data from Insight-HXMT.
+With reflection spectroscopy, these authors found the disk inclination to be $\sim 87^{\circ}$.
+However, none of the studies in the literature discussed the origin of the spectral and temporal properties in the context of the shock oscillation model, which prompted us to analyze the broadband X-ray data from Insight-HXMT and interpret them in light of accretion with shocks.
 The paper is organized as follows: in §2, we describe the observation, data selection, reduction, and analysis procedures.
@@
 The details of the selected data are given in Table 1.
+Here, we note that, in general, the exposure is the time span of observation within an observation ID.
+The Insight-HXMT has several observation IDs for this source, each of which lasts for a different amount of time, which is the exposure time.
+In Table 1, one can see that all the observation IDs have different exposures, with some of them having a time gap between them.
+When reducing the data, we found that inside every observation ID, there were several sub-IDs, the name of which starts with that particular ID.
+The total exposure of an entire observation ID was cut into several sub-IDs with different exposure times but with no time gap between them.
+This can be seen in Table 3 (columns 3 and 5).
+GTIs are also different.
+They can be separated in time between the consecutive ones.
+Here, this is not the case.
+Thus, to differentiate between them, we called these as ‘exposure IDs’, and used this phrase throughout the paper.
+Within 1 observation ID (like the first one in Table 1), there are several exposure IDs (like 12 exposure IDs in the 1st column of Table 3).
+This has not been defined like this before.
+To keep things easy, we have taken this approach.
 Table 1: List of observations used for this work.
@@
 (1) (2) (3) (4) (5) (6)
-P0614338001 2023-08-25 2023-08-27 60181.3410 60183.0597 148495
-P0614338002 2023-08-27 2023-08-28 60183.0597 60184.2327 137343
-P0614338003 2023-08-29 2023-08-31 60185.3075 60187.1589 159958
-P0614338004 2023-08-31 2023-09-02 60187.1588 60189.0765 165684
-P0614338005 2023-09-02 2023-09-04 60189.0765 60191.0601 171384
-P0614338006 2023-09-04 2023-09-06 60191.0601 60193.4447 206034
-P0614338008 2023-09-07 2023-09-08 60194.0350 60195.0926 91378
-P0614338009 2023-09-08 2023-09-09 60195.0926 60196.0840 85659
-P0614338010 2023-09-09 2023-09-10 60196.0840 60197.0754 85653
-P0614338011 2023-09-10 2023-09-11 60197.0753 60198.0667 85650
+P0614338001 2023-08-25 2023-08-27 60181.34 60183.05 148495
+P0614338002 2023-08-27 2023-08-28 60183.05 60184.23 137343
+P0614338003 2023-08-29 2023-08-31 60185.30 60187.15 159958
+P0614338004 2023-08-31 2023-09-02 60187.15 60189.07 165684
+P0614338005 2023-09-02 2023-09-04 60189.07 60191.06 171384
+P0614338006 2023-09-04 2023-09-06 60191.06 60193.44 206034
+P0614338008 2023-09-07 2023-09-08 60194.03 60195.09 91378
+P0614338009 2023-09-08 2023-09-09 60195.09 60196.08 85659
+P0614338010 2023-09-09 2023-09-10 60196.08 60197.07 85653
+P0614338011 2023-09-10 2023-09-11 60197.07 60198.06 85650
 2.2 Data Reduction
@@
 The PDS is normalized in a way that their integral can give the RMS-squared fractional variability.
-The white noise level is subtracted and a geometrical rebinning of -1.02 is used.
+The white noise level is subtracted by using a normalization value of $-2$ in powspec and a geometrical rebinning of -1.02 is used.
 For some data, we needed to use a geometrical rebinning of -1.05 to have a prominent QPO-like nature in the PDS.
-We fit the QPO using single or multiple Lorentzian model components in powspec to extract QPO properties like frequency ($\nu_{qpo}$), full-width at half maximum or FWHM, and normalization.
+We model the PDS with Lorentzian model for both the fundamental QPO and the harmonic components in powspec.
+The parameters from the fitted Lorentzian model for the fundamental QPO help us obtain the frequency ($\nu_{qpo}$), the full-width at half maximum or FWHM, and the normalization.
 For several observations, we have found the presence of single or multiple harmonics.
@@
 The spectral analysis was done for the exposures are marked with ‘*’ in Table 3.
-We did not do the spectral analysis for all epochs as spectral variability was not present within such a small timescale (intraday) like the variability in the light curves.
+We did not perform the spectral analysis for all epochs as we did not see change in spectral features over such a small time gap unlike the timing properties, which showed variations within a day.
 We report them in the result section.
@@
 3 Results
-Using the timing and spectral analysis, we have studied the accretion flow properties of the very recent 2023 outburst of the BHC Swift J1727.8-1613.
+Using the timing and spectral analysis presented above, we have studied the accretion flow properties of the very recent 2023 outburst of the BHC Swift J1727.8-1613.
 We studied the QPO properties of the source during the outburst as well as the spectral nature and radiation properties using Insight-HXMT data.
@@
 3.1 Temporal Properties
-Here, we discuss the evolution of the flux and hardness ratio, and QPO properties during the first 10 observation IDs of the outburst.
 3.1.1 Outburst Profile and Hardness Ratio
@@
 In panel (c), we show the variation of the hardness ratio, which is the ratio of the hard ($F_{H}$) $6-20$ keV to the soft ($F_{S}$) $2-6$ keV flux.
-At the start of the outburst, the HR was $\sim 0.5$ and then it gradually decreased to MJD 60235 as the outburst progressed.
+At the start of the outburst, the HR was $\sim 0.5$ and then it gradually decreased till MJD 60235 as the outburst progressed.
 After that, the HR became almost constant, although decreasing very slowly to $\sim 0.025$ at the end of the considered data of the outburst on MJD 60350.
@@
 Without studying the detailed spectral analysis, HR provides a quick idea about the spectral nature of an outburst.
-However, for this source, we notice that HR was a little higher at the onset of the outburst, where the hard flux was also comparable to the soft flux.
-However, after a few days, soft flux completely took over the hard flux, and the HR was also not high.
-Depending on the HR value, we propose that the source was already in the SIMS after the onset of the outburst.
+Depending on the HR value, we propose that the source was already in the intermediate state after the onset of the outburst.
 After MJD 60185, it made a transition into the SS.
-Since we are not analyzing the full duration of the outburst, it is not the scope of the paper to study the full spectral classification.
-We only define the spectral state of the source based on the duration of which data is used in this work.
 3.1.2 Quasi Periodic Oscillations (QPOs)
+Figure 2: Model fitted PDS using $0.01$ sec time-binned HE light curve from the observation ID P0614338004 (Exposure ID P061433800401-20230831-01-01).
+Refer to captionRefer to captionRefer to caption
+Figure 3: Evolution of QPO frequencies (Hz) with time.
+Here we have shown three different plots for (a) LE, (b) ME, and (c) HE instruments.
+Panels 1–10 in each of the three of these plots (a, b, c) represent the $\nu_{qpo}$ variation for the 10 observation IDs we used.
 As described in §2, we used the $0.01$ sec time-binned light curves from all three bands (LE, ME, and HE) and produced the PDS to study QPOs.
 We find that QPO nature was present in all the light curves in all three energy bands.
+In Figure 2, we show a model fitted PDS continuum in the $0.1-50$ Hz frequency range for $0.01$ sec time-binned HE light curve from the observation ID P0614338004 (Exposure ID P061433800401-20230831-01-01).
 In Figure 3, we show the variation of the QPO frequency ($\nu_{qpo}$) for the 10 different obs.
@@
 We find that for every single obs ID, the $\nu_{qpo}$ showed significant variation.
-Figure 2(a), (b), (c) show the variations of the $\nu_{qpo}$ with time for 10 different obs IDs for LE, ME, and HE respectively.
-Evolution of QPO frequency in different observation ID.s
-Refer to captionRefer to captionRefer to caption
-Figure 2: Evolution of QPO frequencies (Hz) with time.
-Here we have shown three different plots for (a) LE, (b) ME, and (c) HE instruments.
-Panels 1–10 in each of the three of these plots (a, b, c) represent the $\nu_{qpo}$ variation for the 10 observation IDs we used.
+Figure 3(a-c) shows the variations of the $\nu_{qpo}$ with time for 10 different obs IDs for LE, ME, and HE respectively.
+It is interesting to observe that $\nu_{qpo}$ has changed in frequency so much within this short period, which suggests that there are some changes happening in the system in a very short timescale.
 Figure 4: Evolution of QPO frequency with time during the whole period of analysis for (a) LE, (b) ME, and (c) HE.
-In Figure 3, we show the evolution of the $\nu_{qpo}$ for the total duration of our analysis period.
+In Figure 4, we show the evolution of $\nu_{qpo}$ for the total duration of our analysis period.
 In (a), (b), and (c) panels the variations are for LE, ME, and HE respectively.
-We find that the QPO frequencies were $\sim 0.23\pm 0.008$, $0.24\pm 0.002$, and $0.24\pm 0.002$ Hz for LE, ME, and HE on the starting day of our analysis.
-After this day, the $\nu_{qpo}$ shows a rapid increase with time till $\sim$ MJD 60187 when $\nu_{qpo}$ was $\sim 1.2$Hz for all the three bands.
-After this date, it increased very slowly till MJD $\sim 60190$ when $\nu_{qpo}$ was $\sim 1.45$.
-Then it decreased for a very short period until it rapidly increased to $1.5$Hz again in all three bands.
-After that, it showed varying nature till MJD $\sim 60197$, and then it finally increased again.
-We have discussed the possible reason behind this nature in the next section.
-Using the PDS fitting, we extracted some information (e.g., full-width at half maximum (FWHM), Normalization (LN)) of QPOs.
-We also extracted source and background count rates for the exposures.
-Using these estimates, we calculated the $Q$-value and RMS that represent the sharpness of the QPO and the fractional variability in the PDS, respectively.
+We find that the QPO frequencies were $0.23\pm 0.01$, $0.24\pm 0.01$, and $0.24\pm 0.01$ Hz for LE, ME, and HE on the starting day of our analysis.
+After this day, $\nu_{qpo}$ shows a rapid increase with time till $\sim$ MJD 60187 when $\nu_{qpo}$ was $1.16\pm 0.01$, $1.16\pm 0.01$, and $1.16\pm 0.01$ for LE, ME, and HE, respectively.
+After this date, it increased very slowly till MJD $\sim 60190$ when $\nu_{qpo}$ was $1.44\pm 0.01$, $1.44\pm 0.01$, and $1.50\pm 0.01$ for the three bands.
+Then it decreased for a very short period until it rapidly increased to $\sim 1.5$ Hz again in all three bands.
+After that, it increased and decreased in a narrow range of $\sim 1.0-1.5$ Hz till MJD $\sim 60197$, and then it finally increased again.
+We have discussed the possible reasons behind this nature in the next section.
+Using the PDS fitting, we extracted the centroid frequency, the full-width at half maximum (FWHM) and the Normalization (LN)) of the QPOs.
+The ‘fplot’ task of the ‘ftools’ package in the HEASoft software produces plots of the light curves.
+Using this task, we plotted and saved each value of time and the corresponding count rate.
+Then we average the counts to find out the average count rate for that light curve.
+The same task is applied to extract both the source and background count rates for all the exposures.
+Then, using these estimates, we calculated the $Q$-value and RMS that represent the sharpness of the QPO and the fractional variability in the PDS, respectively.
 These values are listed in Table 4 in columns 5–7 ($Q$-value) and 8–10 (RMS) for LE, ME, and HE respectively.
-Evolution of RMS (%) with QPO frequency and time
-Refer to captionRefer to caption
-Figure 4: 3-D representation of the variation of $\nu_{qpo}$ and (%)RMS with time (MJD) for (a) LE, (b) ME, and (c) HE.
-In Figure 4, we show a 3-D representation of the evolution of QPO and RMS with time.
+Figure 5: Evolution of the QPO RMS (%) with time (MJD) for (a) LE, (b) ME, and (c) HE.
+In Figure 5, we show the evolution of the QPO RMS (%) with time (MJD).
 In the LE band (Fig.
-4a), the RMS was higher at the start of the analysis period and it then gradually decreased.
+5a), the RMS was higher at the start of the analysis period and it then gradually decreased.
 In the ME band (Fig.
-4b), it increased from the onset of the analysis period for some time and then again decreased.
+5b), it increased from the onset of the analysis period for some time and then decreased.
 The HE band (Fig.
-4c) shows a similar trend as in LE band.
+5c) shows a similar trend as in LE band.
 Since the QPO nature has already been reported in the recent paper by Yu et al.
-(2024), we are not focusing on the classification of the QPO nature.
+(2024), we are not focusing on the classification of the QPO.
 However, by the values of both $Q$-factor and RMS, it can be said that the QPO is type-C in nature for all the exposures reported here.
 3.2 Spectral Properties
+Refer to captionRefer to caption
+Figure 6: Model fitted spectrally analyzed unfolded spectra.
+(a) is or observation ID.
+P061433800501 (Exposure: P061433800501-20230902-01-01) for which the gabs model was required to achieve the best fit.
+(b) is for the observation ID.
+P061433800601 (exposure: P061433800601-20230904-01-01) for which the gabs model was not required to achieve the best fit.
 Along with the timing properties, studying the spectral properties provides more insights on the nature of the outburst.
@@
 We have simultaneously used LE+ME+HE in the $2-150$ keV energy band (LE in 2–10, ME in 10–35, and HE in 35–150 keV) for our spectral fitting for all the selected exposure IDs.
-Model Fitted Insight-HXMT Spectra
-Refer to captionRefer to caption
-Figure 5: Model fitted spectrally analyzed unfolded spectra.
-(a) is or observation ID.
-P061433800501 (Exposure: P061433800501-20230902-01-01) for which the gabs model was required to achieve the best fit.
-(b) is for the observation ID.
-P061433800601 (exposure: P061433800601-20230904-01-01) for which the gabs model was not required to achieve the best fit.
-We first fitted the data with phenomenological diskbb and power-law models.
+We first fitted the data with the phenomenological diskbb and power-law models.
 We used the multiplicative tbabs model to account for the interstellar absorption.
@@
 Thus, we refitted the data using the model combination as constant*tbabs*gabs(diskbb + power-law + ga + pexrav).
-Here, the gabs model takes care of the low energy absorption feature, while the ga and pexrav models take care of the contribution from the Gaussian and hump-like nature, as mentioned above.
+Here, the gabs model takes care of the low energy absorption feature, while the gaussian and pexrav models take care of the contribution from the Gaussian and hump-like nature, as mentioned above.
 In Fig.
 5a, we show the best-fitted unfolded spectrum using this model combination for the observation ID P061433800501 (Exposure ID: P061433800501-20230902-01-01), where $\chi^{2}/DOF\sim 1$.
-After some exposure, we found that the gabs model was not needed anymore to achieve the best fit.
+After some exposures, we found that the gabs model was not needed anymore to achieve the best fit.
 Then, the model combination in XSPEC reads as constant*tbabs(diskbb + power-law + ga + pexrav).
@@
 P061433800601 (exposure ID: P061433800601-20230904-01-01), for which $\chi^{2}/DOF\sim 0.9$.
-Model Fitted Parameters
 Refer to captionRefer to caption
-Figure 6: Variation of spectrally fitted properties (a) hydrogen column density ($n_{H}$), (b) inner-disk temperature ($T_{in}$), (c) diskbb normalization, (d) Photon Index ($\Gamma$), (e) reflection fraction (Rref), and (f) cutoff energy (Ecut) with time.
+Figure 7: Variation of spectrally fitted properties (a) hydrogen column density ($n_{H}$), (b) inner-disk temperature ($T_{in}$ in keV), (c) diskbb normalization, (d) Photon Index ($\Gamma$), (e) reflection fraction (Rref), and (f) cutoff energy (Ecut) with time.
 […]
 3.3 Correlation between spectral and timing properties
-Photon Index ($\Gamma$) vs QPO frequency ($\nu_{qpo}$)
 Refer to captionRefer to captionRefer to caption
@@
 Here, $\Gamma$ is the spectral fitted index using combined LE+ME+HE data.
-The QPO frequency can be produced from the oscillation of the Compton corona, which can also be the origin of the hard power-law photons that extend the spectrum to higher energies.
-Therefore, these two features might be interlinked or correlated.
 It has been previously seen that the QPO frequency ($\nu_{qpo}$) and $\Gamma$ show a positive correlation for some of the sources during their active outbursting phase (Vignarca et al.
@@
 Considering these points, we have found a positive correlation between $\Gamma$ and $\nu_{qpo}$.
-Using both the Pearson-Linear and Spearman-Rank correlation methods, we found correlation coefficients for the three energy bands as shown in the table below.
+Using both the Pearson-Linear and Spearman-Rank correlation methods, we found correlation coefficients for the three energy bands as shown in Table 2.
+In all of the three energy bands, we see moderately strong correlation when it comes to $\Gamma$ and $\nu_{qpo}$ as correlation coefficient is $\sim 0.5-0.8$.
+This shows the timing and spectral properties, the origin of which are the same geometric variation of the Compton corona, are moderately correlated in this source.
+Table 2: Correlation coefficients between the QPO frequency ($\nu_{qpo}$) and photon index of power-law ($\Gamma$).
 Pearson-Linear Spearman-Rank
@@
 HE $0.64\pm 0.15$ $0.64\pm 0.19$
-4 Discussions
+3.4 Propagation of Shock
+Figure 9: Variation of shock properties with time in accordance with $\nu_{qpo}$.
+In (a) the variation of $\nu_{qpo}$ with time and in (b) and (c) we show the variations of the shock location ($X_{s}$), and shock speed ($v$) with time.
+In the TCAF paradigm, the oscillation of the shock is responsible for the origin of QPOs.
+When this shock front oscillates, it produces QPOs.
+The shock oscillation model has two types of timescales: the infall timescale ($t_{i}$) and the cooling timescale ($t_{c}$) and matching of these timescales gives the resonance condition responsible for the QPOs.
+The relation between the QPO frequency and shock location is given as (Molteni et al.
+1996; Chakrabarti & Manickam 2000; Chakrabarti et al.
+2004),
+where, $c$, $G$, $M_{BH}$, $X_{s}$, and $R$ are the speed of light, the gravitational constant, the mass of the BH, location of the shock (in units of the Schwarzschild radius $r_{s}$), and the ratio of matter densities in post-shock to pre-shock regions ($\rho_{+}/\rho_{-}$) respectively.
+From our timing analysis, we extracted the information about the QPO frequency ($\nu_{qpo}$).
+Using the above relation, we estimated the shock location during the outburst.
+Using Eq.
+1, we found that the shock was far away ($>500~{}r_{s},r_{s}=2GM_{BH}/c^{2}$) from the BH at the start of the outburst (see Figure 9b).
+It then shifted towards the black hole sharply as the QPO frequency increased rapidly during the initial few days ($\sim 250~{}r_{s}$).
+Close to MJD 60185, the shock location started to decrease very slowly and it reached to $\sim 100~{}r_{s}$ at the end of our analysis period.
+The values of the shock location are given in Table 5.
+If the time of the first exposure ID was taken to be 0, and the shock location as $X_{s0}$ and after time $t$, if the shock is at $X_{s}$, then the velocity of the movement of the shock would be given by the equation (Chakrabarti et al.
+2005)
+[…]
+Using this relation, we have estimated the velocity of the movement of the shock.
+The estimated shock velocities are given in Figure 9c, which is a few tens of m/s in agreement with the previous estimates (Chakrabarti et al.
+2005; Mondal et al.
+2015).
+We find that the shock velocity varies within this period.
+In the beginning, the shock moved with a speed of $\sim 50-100$ m s-1, when the $X_{s}$ was also larger.
+Then, after MJD $\sim$ 60183.7, it decreased a little bit with values $<50$ m s-1.
+Then it slowed down and was moving with a speed of $\sim 10-30$ m s-1 with some fluctuations and continued till the end of the observation period.
+4 Discussion
 The Galactic black hole Swift J1727.8-1613 went through an outburst recently in 2023.
@@
 Using the spectra files from these three instruments, we also studied the spectral properties of this source by analyzing combined LE+ME+HE spectra in the $2-150$ keV broad energy range.
-4.1 Propagation of Shock
-Figure 8: Variation of shock properties with time in accordance with $\nu_{qpo}$.
-In (a) the variation of $\nu_{qpo}$ with time and in (b) and (c) we show the variations of the shock location ($X_{s}$), and shock speed ($v$) with time.
 Quasi-periodic oscillation is one of the most important and common phenomena for stellar mass black holes.
@@
 It showed an increasing trend from the start of our analysis.
-From the variation of QPO frequency, (%)RMS, and $Q$-factor, the nature of the QPO is designated to be type-C.
-Although the QPOs are extensively studied in the literature, their origin requires in depth modeling.
-Here, we would like to focus on the physical scenario that explains the origin of QPOs due to the shock instabilities in advective flows around BHs.
-Accretion onto BH is transonic flow with possibility of multiple sonic points (Chakrabarti 1989).
+We found that the RMS (%) was very high ($>13\%$) for all the exposures.
+The $Q$-value was also high, though for some observations it was close to those QPOs that are types-A/B.
+However, the frequency and rms allow us to classify the QPOs as type-C, which are in agreement with Casella et al.
+(2005).
+Considering the values of RMS, $Q$-value and the type of QPO, we suggest that the source was in the HIMS.
+Recently, Yu et al.
+(2024) also reported the timing properties of this source for a longer period.
+The QPO frequency has varied in a much broader range in their report from 0.1 to 8 Hz.
+The authors also concluded the observed QPOs in the full period as type-C.
+During the 2009-10 outburst of the BHC XTE J1752-223 (Chatterjee et al.
+2020), we found the existence of type-C QPOs that had $Q$-values comparable to this source.
+During the 2016-17 outburst of the BHC GRS 1716-249 (Chatterjee et al.
+2021), type-C QPOs were found to be present with similar RMS (%) values.
+Other than these sources, Belloni et al.
+(2002), Casella et al.
+(2005), and the references therein can be seen for an analogy.
+Such similarities in observed QPO properties in different sources require physical explanations of common origin.
+Here, we would like to focus on the physical scenario that explains the origin of QPOs due to the shock oscillations in advective flows around BHs.
+Accretion onto BH can be explained as transonic flow with possibility of multiple sonic points (Chakrabarti 1989).
 The matter from the companion need not necessarily be only Keplerian.
@@
 Additionally, it has also been observed by the authors that once the QPO sets in, it gets locked for a week or more depending on the above condition.
-From the expressions of infall and cooling timescales, we get that the QPO frequency ($\nu_{qpo}$) can be estimated using the relation (Molteni et al.
-1996; Chakrabarti & Manickam 2000; Chakrabarti et al.
-2004),
-Where $c$, $G$, $M_{BH}$, $X_{s}$, and $R$ are the speed of light, gravitational constant, the mass of the BH, location of the shock, and the ratio of matter densities in post-shock to pre-shock regions ($\rho_{+}/\rho_{-}$) respectively.
-From our timing analysis, we already extracted the information about QPO frequency ($\nu_{qpo}$).
-Using the above relation, we estimated the shock location during the outburst.
 According to Chakrabarti et al.
@@
 2015).
-Using Eq.
-1, we found that the shock was far away from the BH at the start of the outburst (see Figure 8b).
-It then significantly decreased as the QPO frequency increased rapidly during the initial few days.
-Close to MJD 60185, when the $\nu_{qpo}$ stopped increasing, and the shock location started to decrease very slowly.
-The values of the shock location along with QPO frequency are given in Table 3.
-If the time of the first exposure ID was taken to be 0, and the shock location as $X_{s0}$ and after time $t$, if the shock is at $X_{s}$, then the velocity of the movement of the shock would be given by the equation (Chakrabarti et al.
-2005)
-[…]
-Using this relation, we have estimated the velocity of the movement of the shock.
-The estimated shock velocities are given in Figure 8c, which is a few tens of m/s in accord with the previous estimations (Chakrabarti et al.
-2005; Mondal et al.
-2015).
-We find that the shock velocity has varied within this period.
-At the beginning, it attained higher values when the $X_{s}$ was also larger, however decreasing significantly afterward.
-It slowed down as the shock moved inwards, along with some fluctuations.
-This can be due to the limitation of assuming constant velocity of the shock propagation.
-4.2 Evolution of the physical configuration and spectral nature
+When the cooling timescale ($t_{c}$) and infall timescale ($t_{i}$) becomes comparable to each other (Molteni et al.
+1996), an oscillation could set in the shock boundary due to resonance.
+This could produce variations as we observe in the light curves.
+This might be the case for this source regarding the origin of the occurrence of the QPOs.
+Here, from this analysis, we have found that the shock was far away ($\sim 500~{}r_{s}$) when QPO frequency ($\nu_{qpo}$) was low ($\sim 0.2$ Hz).
+The disk was far away.
+As the disk moves inward, cooling increases with progressive days and QPO frequency increases, as shown in Eq.
+1; following inverse scaling with $X_{s}$.
+This is the case that we observed here.
+After the start, as the outburst progressed, the cooling began to take place and $X_{s}$ got smaller in size.
+As a result, $\nu_{qpo}$ increased.
+After a few days, the shock location decreased very slowly, without changing its value much, the $\nu_{qpo}$ was also increasing very slowly.
+We also tried to observe how the shock moved, considering the constant velocity approach, as given in Eq.
+2.
+We have found that the shock speed varied in a broad range.
+At the start, the shock velocity was relatively higher compared to later days, while it became nearly constant with some fluctuations in later time.
+Such fluctuations can be due to the limitation of assuming constant velocity of the shock propagation.
+For a comparison, we refer to Chakrabarti et al.
+(2005).
+The authors showed that considering the constant velocity approximation, the average speed of the shock is close to 20 m s-1, which is nearly consistent with our result.
+This also depends on the non-constant time difference between the exposures.
+However, the overall velocity profile agrees that the shock moved faster when $X_{s}$ was bigger, decreased with $X_{s}$, and afterward almost with constant velocity.
 From the spectral properties, we have seen that the $T_{in}$ was low initially and then gradually increased.
@@
 Furthermore, this also supports the properties we showed from the timing analysis that the shrinking of post-shock region with time.
-Looking at the spectral parameters, and also the location of the shock, we designate that the source was already in an intermediate state at the start of our analysis period.
+The photon index of power-law ($\Gamma$) is an important property to decide the spectral nature of an ongoing outburst.
+This decides the slope in the spectrum, which tells whether the spectrum extends up to high energy or not.
+When an outburst stays in the hard state, the photon index of power-law ($\Gamma$) generally stays low $\sim 1.7$ (Remillard & McClintock 2006).
+From the expression of power-law component ($A(E)\propto E^{-\Gamma}$), this suggests that the spectrum extends up to high energy.
+As the source gets towards intermediate states, $\Gamma$ becomes $\sim 2$ (e.g., Remillard & McClintock 2006), even up to 3, sometimes.
+In the high soft state, $\Gamma$ reaches a value of $>3$ (Remillard & McClintock (2006), for a review), although sources have been discovered to reach soft state with $\Gamma<3$.
+According to the TCAF configuration (Chakrabarti & Titarchuk 1995), as discussed earlier in the Introduction, the shock location, that produces both the change in the spectral and timing properties, changes in size in different spectral states.
+In the HS it stays very far from the source, at $\sim 1000~{}r_{s}$.
+The disk is even further away.
+As a result, only a few photons come in the process of radiation.
+As the disk approaches closer, the $X_{s}$ becomes smaller and the component from the disk starts to gradually dominate over the hard power-law component.
+In the intermediate states, these two components stay comparable.
+This time the shock stays at a few $100~{}r_{s}$.
+In the SS, the disk moves even closer and $X_{S}$ becomes very small (a few tens of $r_{s}$).
+Thus, the disk component now fully dominates, and the spectrum becomes very soft.
+Thereby, the location of the shock determines the spectral state of the outburst.
+Looking at the spectral parameters, and also the location of the shock, we propose that the source was already in an intermediate state at the start of our analysis period.
 As the outburst progressed, it was transitioning towards softer states.
 From the values of the parameter $\cos(i)$ of the pexrav model, we find that the inclination of the disk was $>60^{\circ}$, which makes this source a highly inclined system.
-The hydrogen column density ($N_{H}$) showed a variation, which is common for Galactic black hole binaries unless the system is heavily obscured, which has recently been found for the newly discovered BH Swift J151857.0-572147 (Mondal et al.
-2024).
-The average $N_{H}$ has an average value of $0.27_{-0.17}^{+0.08}\times 10^{22}$ cm-2.
+This is in agreement with the recent reports of Bouchet et al.
+(2024), Yu et al.
+(2024).
+The hydrogen column density ($N_{H}$) showed a variation, which is common for Galactic black hole binaries (Eckersall et al.
+2017; Chatterjee et al.
+2023) unless the system is heavily obscured, which has recently been found for the newly discovered BH Swift J151857.0-572147 (Mondal et al.
+2024), Swift J1658.2-4242 (Mondal & Jithesh 2023 and references therein).
+The $N_{H}$ has a variation in the range of $(0.12\pm 0.02-0.39\pm 0.08)\times 10^{22}$ cm-2.
+The QPO frequency can be produced from the oscillation of the Compton corona, which can also be the origin of the hard power-law photons that extend the spectrum to higher energies (e.g., Lee & Miller 1998; Kumar & Mishra 2014; Karpouzas et al.
+2020; Liu et al.
+2022; Bellavita et al.
+2022; etc.).
+Therefore, these two features might be interlinked or correlated.
+As described in the introduction, the oscillation of the CENBOL can be the mechanism producing the QPOs.
+This CENBOL is also responsible for producing the power-law part of the spectrum.
+The change in the size of the shock location gives the change of spectral states as well as the frequency of the QPOs.
+We found such correlations for the source Swift J1727.8-1613 during its 2023 outburst.
 5 Summary and Conclusions
@@
 From our analysis, we conclude that:
-(i) Quasi-periodic oscillations were present during the entire period of our analysis, which showed evolution in QPO frequency.
-(ii) The $Q$-factor and $RMS$ help designate the nature of the QPOs as type-C.
-(iii) QPOs originated from the shock instabilities that was formed by matter having an angular momentum distribution deviated from the Keplerian distribution.
-As the shock moved inwards, the QPO frequency was observed to increase.
+(i) Quasi-periodic oscillations were present during the entire period of our analysis, which showed evolution in QPO frequency from $0.21\pm 0.01$ to $1.86\pm 0.01$ Hz within this analysis period.
+(ii) The values of the $Q$-factor and $RMS$ best represent the QPOs as type-C.
+(iii) Our study infers that the QPOs might have originated due to the resonance oscillations of the shocks. As the shock moved inwards (outwards), the QPO frequency increased (decreased).
 (iv) With the shock moving inwards, the spectral nature of the source was transitioning towards the softer states, after it started in the intermediate state at the onset of the analysis period.
@@
 (vi) This source has a high inclination.
-(vii) The hydrogen column density varied in the range of $0.10-0.35\times 10^{22}$ cm-2, which is quite common for interstellar absorption of Galactic black holes.
+(vii) The hydrogen column density varied in the range of $(0.12\pm 0.02-0.39\pm 0.08)\times 10^{22}$ cm-2., which is quite common for interstellar absorption of Galactic black holes.
 6 Data Availability
@@
 7 Acknowledgements
+We thank the scientific editor and the anonymous referee(s) for their detailed comments and insightful suggestions that improved the quality of the paper.
 KC acknowledges support from the SWIFAR postdoctoral fund of Yunnan University.
-SM acknowledges the Ramanujan Fellowship (# RJF/2020/000113) by SERB-DST, Govt.
+SM acknowledges the Ramanujan Fellowship (# RJF/2020/000113) by SERB/ANRF-DST, Govt.
 of India for this research.